### v4

# Triangle processes on graphs with given degree sequence

Abstract
The switch chain is a well-studied Markov chain which generates random graphs with a given degree sequence and has uniform stationary distribution. Motivated by the high number of triangles seen in some real-world networks, we study a variant of the switch chain which is more likely to produce graphs with higher numbers of triangles. Specifically, we apply a Metropolis scheme designed to have the following stationary distribution: graph $G$ has probability proportional to $\lambda^{\min\{t(G),\nu\}}$, where $t(G)$ is the number of triangles in $G$ and $\nu$ is a cut-off value introduced to moderate the impact of graphs with a very high number of triangles. We assume that the “activity” $\lambda$ satisfies $\lambda\geq 1$, and call the resulting chain the modified Metropolis switch chain. We prove that the modified Metropolis switch chain is rapidly mixing whenever the (standard) switch chain is rapidly mixing, provided that the activity and maximum degree are not too large.
The triangle switch (or $\triangle$-switch”) chain is a restriction of the switch chain which only performs switches that change the set of triangles in the graph. We prove that the $\triangle$-switch chain is irreducible for any degree sequence with minimum degree at least 3, and prove a rapid mixing result for the modified Metropolis $\triangle$-switch chain.
Finally, we investigate the distribution of triangles in random graphs with given degrees, under both the uniform distribution and the distribution in which graph $G$ has probability proportional to $\lambda^{t(G)}$. Our analysis implies that the imposition of the cut-off $\nu$ does not significantly impact the behaviour of these modified Metropolis chains over polynomially many steps.
1 Introduction
Randomly generating graphs has been the subject of considerable research interest. See, for example, [1, 2, 3, 9, 14, 17, 22, 27, 44, 52, 47, 61]. A comprehensive survey of work to date is given in [31].
Generation using Markov chains has been an important approach in this area, particularly Markov chains based on switches, for example [3, 14, 17, 22, 44, 46, 52, 47, 61]. Switches delete a pair of edges from the graph and insert a pair using the same four vertices. They have the important property that they preserve the degree sequence of the graph. Thus they are useful for generating regular graphs, or other graphs with a given degree sequence. Markov chains also give a dynamic reconfigurability property, which is useful in some applications, for example [14, 22, 47].
For any such Markov chain, three questions arise. Firstly, can the chain generate any graph in the chosen class? In other words, is the Markov chain ergodic? Secondly, if so what is its equilibrium distribution? Thirdly, what is the rate of convergence to equilibrium? That is, what is the mixing time of the chain? Note that it is usually fairly easy to ensure that a Markov chain is aperiodic, so the second question focusses on establishing irreducibility.
The expected number of triangles in a uniformly random graph with a given degree sequence can be bounded above by the cube of the maximum degree. Hence, for example, when all degrees are bounded we expect only a bounded number of triangles in this random graph. However, many real-world networks, such as social networks, contain many triangles [29, 43]. This motivated us to adapt the transition procedure of the switch chain by introducing a Metropolis accept/reject scheme [36, 51], described in Section 1.2. This chain has an equilibrium distribution $\hat{\pi}_{\lambda}$ in which the probability of a graph is exponentially weighted by its number of triangles, up to some cut-off which is required to moderate the impact of graphs with a very high number of triangles. The chain has a parameter $\lambda\geq 1$ which is a fixed real number. Our first result (Theorem 1.1) shows that this chain, which we call the modified Metropolis switch chain, is rapidly mixing whenever the (standard) switch chain is rapidly mixing, under some conditions on the degree sequence and the parameter $\lambda$ of the chain.
The triangle switch Markov chain, which we abbreviate as $\triangle$-switch chain”, was introduced in [15] to explore ways to generate regular random graphs with more triangles than would appear under the uniform distribution. See Section 1.2 below for the relevant definitions. Transitions of the $\triangle$-switch chain are precisely those switches which change the set of triangles in the graph; several options for the transition probabilities were proposed in [15]. These transition probabilities were designed to encourage the formation of more triangles in fewer steps, compared with the standard switch chain. One variant was shown to produce $\Omega(n)$ triangles in $O(n)$ steps of the chain [15, Theorem 2], though it is not known whether that variant of the chain is close to stationarity after $O(n)$ steps.
The $\triangle$-switch chain was proved to be irreducible for 3-regular graphs (note, irreducibility depends only on the set of transitions with positive probability, and not the values of the transition probabilities). The proof of irreducibility involved finding a sequence of transitions of the chain ($\triangle$-switches) which transform any 3-regular graph into disjoint copies of $K_{4}$ together with a disjoint 6-vertex graph, and then showing that any such graph can be transformed into another using $\triangle$-switches. This proof is not well suited to standard methods for proving rapid mixing of the chain. A further obstacle is that some of the variants studied in [15] do not seem to be time-reversible, a requirement for most approaches to proving rapid mixing.
The proof of irreducibility in [15] was extended in [16] to $d$-regular graphs, for any $d$. The proof involved finding a sequence of $\triangle$-switches which transform any $d$-regular graph into disjoint copies of $K_{d+1}$ together with a disjoint graph, called a “fragment”, with at most $2d$ vertices; then showing that any such graph can be transformed into another using $\triangle$-switches. This required a different approach to proving irreducibility of the fragments, by showing that any step of the flip chain of [47] can be simulated by at most three $\triangle$-switches.
Our second result (Theorem 1.2) extends and generalises the result of [15] by showing that any step of the switch chain [14] can be simulated by at most five $\triangle$-switches. Moreover this holds for all graphs with a given degree sequence such that the minimum degree is at least 3.
Next we apply the Metropolis accept/reject approach to the $\triangle$-switch chain, producing a Markov chain which we called the modified Metropolis $\triangle$-switch chain. This chain has the same stationary distribution $\hat{\pi}_{\lambda}$ as the modified Metropolis switch chain. Our third result (Theorem 1.3) analyses the mixing time of the modified Metropolis $\triangle$-switch chain, using the simulation paths that we introduced to prove irreducibility.
Finally, we analyse the number of triangles in a randomly chosen graph with given degree sequence, under two distributions: the uniform distribution, and the Gibbs distribution $\pi_{\lambda}$ which weights each graph exponentially by the its number of triangles (without any cut-off). This analysis, presented in Section 4, may be of independent interest. We use it to show that the introduction of the cut-off (which limits the impact of graphs with a very high number of triangles) does not noticeably affect the performance of the chain over polynomial-length runtimes.
Our main results are presented formally in Section 1.3, where we also describe the structure of the rest of the paper.
We only know of a few other works which rigorously analyse Markov chains for generating graphs from a known non-uniform distribution. It may be possible that the algorithm of Jerrum, Sinclair and Vigoda [42] for the permanent could be used to sample bipartite graphs with given degree sequence from a non-uniform distribution by adjusting the edge weights, though to our knowledge this has not been explored. Bhamidi, Bresler and Sly [10] analysed the Glauber dynamics (single-edge update) for sampling from the exponential random graph model, and proved rapid mixing in the “high temperature regime”, but also proved that the graphs generated in this regime are very close to binomial random graphs $G(n,p)$, in the sense that any finite set of edges are asymptotically independent. Recently, Jenssen, Perkins, Potukuchi and Simkin [39] analysed the Glauber dynamics for generating from $G(n,p)$ conditioned on being triangle-free, and proved that this Markov chain is rapidly mixing when $p\leq cn^{-1/2}$ for some sufficiently small $c>0$.
Before proceeding, we mention some related work on sampling graphs with a given degree sequence and additional constraints. Mahlmann and Schindelhauer [47] proposed a restricted set of switches, called $k$-flips, which maintain connectivity. The 1-flip chain is rapidly mixing on the set of $d$-regular connected graphs, for any $d$ [17, 22]. In some applications it can be useful to specify the number of edges between sets of vertices with given degrees, as well as the degree sequence. This is done with a joint degree matrix. Work on the switch chain in this context can be found in [4, 18, 21, 58].
1.1 Notation and definitions
Let $[n]$ denote the set $\{1,2,\ldots,n\}$ and let $G=(V,E)$ be an $n$-vertex (simple) graph with vertex set $V=[n]$. The degree sequence $\boldsymbol{d}=(d_{1},d_{2},\ldots,d_{n})$ of $G$ is given by $d_{i}=\deg(i)$ for all $i\in[n]$. An arbitrary sequence $\boldsymbol{d}$ of $n$ nonnegative integers is called graphical if there is any graph which has degree sequence $\boldsymbol{d}$. Given the sequence, this property can be tested efficiently and constructively using, for example, the Havel-Hakimi algorithm [34, 37].
We assume without loss of generality that $d_{1}\geq d_{2}\geq\cdots\geq d_{n}$, so $d_{1}$ is the maximum degree and $d_{n}$ the minimum degree. We note that $d_{1},d_{n}$ are usually denoted by $\Delta,\delta$, but we will not do so here. If $d_{1}=d_{n}=d$, the $d$-regular case, then $d$ is common degree of all vertices. Define
$M=M(\boldsymbol{d})=\sum_{i=1}^{n}d_{i}=2|E|=n\bar{d},\qquad M_{2}=M_{2}(\boldsymbol{d})=\sum_{i=1}^{n}d_{i}(d_{i}-1),$
where $\bar{d}$ is the average degree.
We will denote the number of triangles in $G$ by
$t(G)=\big{|}\big{\{}\{i,j,k\}\subseteq V:ij,jk,ki\in E\big{\}}\big{|}.$
If $G$ has degree sequence $\boldsymbol{d}$ then the maximum value for $t(G)$ is $\nicefrac{{1}}{{3}}\sum_{i=1}^{n}\binom{d_{i}}{2}=M_{2}/6$. (The sum is over all ways of choosing two edges of the triangle meeting a given vertex, the divisor $\nicefrac{{1}}{{3}}$ because this counts each triangle three times.) In the regular case, this maximum is achieved by a $d$-regular graph consisting of the disjoint union of $n/(d+1)$ cliques of size $(d+1)$.
Let $\mathcal{G}_{n,\boldsymbol{d}}$ be the set of all labelled graphs with degree sequence $\boldsymbol{d}$. The switch chain is a Markov chain which walks on $\mathcal{G}_{n,\boldsymbol{d}}$ by choosing two non-incident edges uniformly at random, and randomising these two edges without changing the degree sequence.
It is known that the switch chain converges to the uniform distribution on $\mathcal{G}_{n,\boldsymbol{d}}$, see [23, 46]. Here we consider a also restricted switch chain, in which a switch is only allowed if it changes the set of triangles in the graph. This is achieved by $\triangle$-switches. A $\triangle^{+}$-switch is designed to increase the number of triangles, and a $\triangle^{-}$-switch to decrease this number. See Fig. 1, where $a_{2}a_{1}va_{3}a_{4}$ is a path in $G$ such that $a_{1}a_{3}$ and $a_{2}a_{4}$ are non-edges, which is transformed to a triangle $va_{1}a_{3}$ and a vertex-disjoint edge $a_{2}a_{4}$ and vice versa.
[…]
Fig. 1: The $\triangle$-switches
We denote the transitions in the triangle process as $\triangle$-switches. We may use $\triangle^{+}$-switch to denote a switch which is known to create a triangle, and $\triangle^{-}$-switch for a switch which is known to break a triangle. However, a given $\triangle$-switch can simultaneously create one or more triangles and break one or more triangles, in which case it can be viewed as either a $\triangle^{+}$-switch or a $\triangle^{-}$-switch. We reserve the term “switch”, without qualification, for the transitions of the switch chain, as studied in [14], for example. The fact that a single $\triangle$-switch can destroy and introduce several triangles at once significantly complicates the analysis. Working with a Metropolis chain sidesteps this issue, as it allows us to specify the stationary distribution of the chain.
Let
$\mu(\boldsymbol{d})=M_{2}^{3}/(6M^{3}).$ (1)
We show in Lemma 4.3 below that the expected number of triangles in a uniformly chosen random graph from $\mathcal{G}_{n,\boldsymbol{d}}$ is $\mu(\boldsymbol{d})$ asymptotically if $d_{1}=o(\sqrt{M})$. The Metropolis switch chain introduced in Section 1.2 (see Figure 2) has a non-uniform stationary distribution $\pi_{\lambda}$, where graph $G$ has probability proportional to $\lambda^{t(G)}$, for some parameter $\lambda=\lambda(n)\geq 1$, which we call the activity. When $\lambda>1$ the Metropolis switch chain will generate graphs with more triangles than expected in a uniform random graph with the same degree sequence. In fact, we will show that $\mathbb{E}_{\pi_{\lambda}}[t(G)]\sim\lambda\mu$ in this distribution, if $\lambda$ and $d_{1}$ are small enough. (See Theorem 4.1.)
In order to prove rapid mixing we must modify this chain slightly, to decrease the impact of graphs with an extremely high number of triangles, as described in Section 1.2 (see Figure 4). To demonstrate that this modification does not significantly impact the chain (over a polynomial number of steps), in Section 4 we study the asymptotic distribution of the number of triangles under both the uniform distribution on $\mathcal{G}_{n,\boldsymbol{d}}$ and the stationary distribution $\pi_{\lambda}$ of the Metropolis switch chain. We prove that (other than in the upper tail) the distribution in both cases is asymptotically Poisson, under the assumption that $M_{2}\geq M$. (See Section 4.4 for some discussion about this assumption.) These results may be of independent interest.
There has been a lot of work on the distribution of triangle counts in the Erdős–Renyi model, see for example [5] and references therein. However, prior work on this problem for graphs with given degree sequences was restricted to the regular case, see [24, 28, 50]. In particular, the number of triangles in a random $d$-regular graph is asymptotically Poisson when $d=o(n^{1/5})$, as proved by McKay, Wormald and Wysocka [50]. Z. Gao and Wormald [28] proved that the number of triangles in random $d$-regular graphs is asymptotically normal for $d=o(n^{2/7})$, and this was improved to $d=O(n^{1/2})$ by P. Gao [24].
1.2 Markov chains
For $G\in\mathcal{G}_{n,\boldsymbol{d}}$, let $\mbox{N}(v)=\mbox{N}_{G}(v)$ denote the set of neighbours of $v\in[n]$ and recall that $t(G)$ denotes the number of triangles in $G$.
We now introduce the Metropolis switch chain with parameter $\lambda\geq 1$: the transition matrix of this chain is shown in Figure 2. The standard switch chain corresponds to choosing $\lambda=1$.
At the current state $G\in\mathcal{G}_{n,\boldsymbol{d}}$ do choose a pair of non-adjacent edges $F=\{a_{1}a_{2},\,a_{3}a_{4}\}$ uniformly at random choose a perfect matching $F^{\prime}$ of $\{a_{1},a_{2},a_{3},a_{4}\}$ uniformly at random let $H:=([n],E^{\prime})$ where $E^{\prime}:=\big{(}E(G)\setminus F\big{)}\cup F^{\prime}$ if $F^{\prime}\cap\big{(}E(G)\setminus F\big{)}=\emptyset$ then the next state is $H$ with probability $\min\{1,\lambda^{t(H)-t(G)}\}$ otherwise stay at $G$ end if end
Fig. 2: The transition procedure of the Metropolis switch chain with parameter $\lambda$ on $\mathcal{G}_{n,\boldsymbol{d}}$
Let $P_{\lambda}$ denote the transition matrix of the Metropolis switch chain. At every step we have $P_{\lambda}(G,G)\geq 1/3$, since we choose $F^{\prime}=F$ with probability $1/3$. Hence the Metropolis switch chain is aperiodic, and it is known that all graphs with a given degree sequence are connected by switches [23, 46]. In other words, the Metropolis switch chain is irreducible on $\mathcal{G}_{n,\boldsymbol{d}}$ for all $n$ and $\boldsymbol{d}$.
Define the probability distribution $\pi_{\lambda}$ by $\pi_{\lambda}(G)=\lambda^{t(G)}/Z_{\lambda}(\boldsymbol{d})$ for all $G\in\mathcal{G}_{n,\boldsymbol{d}}$, where
$Z_{\lambda}(\boldsymbol{d})=\sum_{H\in\mathcal{G}_{n,\boldsymbol{d}}}\lambda^{t(H)}$
is the normalising factor. Let $a(\boldsymbol{d})$ be the number of unordered pairs of non-incident distinct edges in any element of $\mathcal{G}_{n,\boldsymbol{d}}$, given by
$a(\boldsymbol{d})=\binom{M/2}{2}-\frac{M_{2}}{2}.$ (2)
If $P_{\lambda}(G,H)>0$ and $G\neq H$ then, assuming without loss of generality that $t(H)\leq t(G)$, we have
$\pi_{\lambda}(G)\,P_{\lambda}(G,H)=\frac{\lambda^{t(G)}}{Z_{\lambda}(\boldsymbol{d})}\,\frac{1}{3a(\boldsymbol{d})}\,\lambda^{t(H)-t(G)}=\frac{\lambda^{t(H)}}{Z_{\lambda}(\boldsymbol{d})}\,\frac{1}{3a(\boldsymbol{d})}=\pi_{\lambda}(H)\,P_{\lambda}(H,G).$
Hence the Metropolis switch chain satisfies the detailed balance equations with respect to $\pi_{\lambda}$, so $\pi_{\lambda}$ is the unique stationary distribution of the chain.
At the current state $G\in\mathcal{G}_{n,\boldsymbol{d}}$ do choose a pair of non-adjacent edges $F=\{a_{1}a_{2},\,a_{3}a_{4}\}$ uniformly at random choose a perfect matching $F^{\prime}$ of $\{a_{1},a_{2},a_{3},a_{4}\}$ uniformly at random let $H:=([n],E^{\prime})$ where $E^{\prime}:=\big{(}E(G)\setminus F\big{)}\cup F^{\prime}$ if $F^{\prime}\cap\big{(}E(G)\setminus F\big{)}=\emptyset$ then the next state is $H$ with probability $\min\{1,\lambda^{\min\{t(H),\nu\}-\min\{t(G),\nu\}}\}$ otherwise stay at $G$ end if end
Fig. 3: The transition procedure of the modified Metropolis switch chain with parameter $\lambda$ on $\mathcal{G}_{n,\boldsymbol{d}}$, where $\nu$ is defined in (3)
For technical reasons, we must also introduce a modified Metropolis switch chain, obtained by replacing $t(H)$ by $\min\{t(H),\nu\}$ and replacing $t(G)$ by $\min\{t(G),\nu\}$ in the transition procedure Figure 2, where
$\nu=\nu(n):=\log n/\log\log n.$ (3)
Note that this does not change the set of transitions, but “dampens” the acceptance probability for states with a very high number of triangles. (Here and throughout the paper, $\log$ denotes the natural logarithm.) The transition procedure of the modified Metropolis chain with parameter $\lambda\geq 1$ is given in Figure 3.
Adapting the detailed balance argument given above shows that the modified Metropolis switch chain has stationary distribution $\hat{\pi}_{\lambda}$ which assigns to graph $G\in\mathcal{G}_{n,\boldsymbol{d}}$ a probability proportional to $\lambda^{\min\{t(G),\nu\}}$. We will see in Section 4 that the modified Metropolis switch chain is polynomial-time indistinguishable from the Metropolis switch chain when $\lambda$ and $\mu$ are small enough: see Remark 4.
We can also apply the same machinery to the triangle switch chain which, recall, only performs transitions that change the set of triangles in the graph. The transition procedure of the modified Metropolis $\triangle$-switch chain is shown in Figure 4. (As above, the word “modified” indicates the use of $\min\{t(G),\nu\}$, rather than $t(G)$.) We prove in Section 2 below that the modified Metropolis $\triangle$-switch chain is irreducible, and hence ergodic, provided the minimum degree $d_{n}\geq 3$. The stationary distribution of the modified Metropolis $\triangle$-switch chain is $\hat{\pi}_{\lambda}$, by detailed balance again.
At the current state $G\in\mathcal{G}_{n,\boldsymbol{d}}$ do choose a pair of non-adjacent edges $F=\{a_{1}a_{2},\,a_{3}a_{4}\}$ uniformly at random choose a perfect matching $F^{\prime}$ of $\{a_{1},a_{2},a_{3},a_{4}\}$ uniformly at random let $H:=([n],E^{\prime})$ where $E^{\prime}:=\big{(}E(G)\setminus F\big{)}\cup F^{\prime}$ if $F^{\prime}\cap\big{(}E(G)\setminus F\big{)}=\emptyset$ and   XXXXXX $\big{(}\mbox{N}(a)\cap\mbox{N}(a^{\prime})\big{)}\setminus\{a_{1},a_{2},a_{3},a_{4}\}\neq\emptyset$ for some $aa^{\prime}\in F\cup F^{\prime}$ then the next state is $H$ with probability $\min\{1,\lambda^{\min\{t(H),\nu\}-\min\{t(G),\nu\}}\}$ otherwise stay at $G$ end if end
Fig. 4: The transition procedure of the modified Metropolis $\triangle$-switch chain with parameter $\lambda$ on $\mathcal{G}_{n,\boldsymbol{d}}$, where $\nu$ is defined in (3)
Comparing the transition procedure in Fig. 4 with the $\triangle$-switch from Fig. 1, we see that taking $aa^{\prime}=a_{1}a_{3}$ satisfies the condition $\big{(}\mbox{N}(a)\cap\mbox{N}(a^{\prime})\big{)}\setminus\{a_{1},a_{2},a_{3},a_{4}\}\neq\emptyset$.
1.3 Our results
Our first result shows that if the activity $\lambda$ and the maximum degree are sufficiently small, then the modified Metropolis switch chain is rapidly mixing whenever the (standard) switch chain is rapidly mixing. The proof uses the “two-stage direct canonical path construction” method from [17].
Theorem 1.1.
Let $\mathcal{D}$ be a family of graphical sequences such that for all $\boldsymbol{d}\in\mathcal{D}$ of length $n$ we have $d_{1}\geq d_{2}\geq\cdots\geq d_{n}\geq 3$. Let $\lambda=\lambda(n)\geq 1$ be a function of $n$. Suppose that there exists an $\alpha\in(0,1)$ such that $\lambda\mu(\boldsymbol{d})\leq\log^{\alpha}n$ for every $\boldsymbol{d}\in\mathcal{D}$ of length $n$ and all large enough $n$. If the switch chain is rapidly mixing on $\mathcal{G}_{n,\boldsymbol{d}}$ for all $\boldsymbol{d}\in\mathcal{D}$ then the same is true for the modified Metropolis switch chain with parameter $\lambda$.
We will use Theorem 1.1 as a convenient stepping-stone in our proof of Theorem 1.3, stated below.
Next we turn to the $\triangle$-switch chain. Recall that a Markov chain is irreducible if the underlying graph of the Markov chain is connected. This graph has an edge corresponding to each transition which occurs with positive probability. Hence, if two chains have the same set of positive-probability transitions and one is irreducible then so is the other.
We show that the $\triangle$-switch chain (or, equivalently, the modified Metropolis $\triangle$-switch chain) is irreducible on $\mathcal{G}_{n,\boldsymbol{d}}$. This was proved for regular graphs in [16], by showing that (a) every $d$-regular graph can be transformed by a sequence of $\triangle$-switches into the disjoint union of $\lfloor n/(d+1)\rfloor$ copies of $K_{d+1}$ and one other small component, called a “fragment”, and (b) using $\triangle$-switches, it is possible to transform any set of $\lfloor n/(d+1)\rfloor$ components, each isomorphic to $K_{d+1}$, into any other such set of components, and to transform fragments into other fragments. This approach is not suitable for proving rapid mixing of the $\triangle$-switch chain by existing methods. Here we will use the “two-stage direct canonical path construction” method from [17] to prove rapid mixing, and this informs our approach to proving irreducibility.
It is known that the set of graphs with a given degree sequence is connected under switches [23, 46]. In other words, the switch chain is irreducible on $\mathcal{G}_{n,\boldsymbol{d}}$ for all $n$ and $\boldsymbol{d}$. Hence in order to prove that the $\triangle$-switch chain is irreducible on $\mathcal{G}_{n,\boldsymbol{d}}$, it suffices to show that every switch $(G,H)$ can be simulated using a sequence of $\triangle$-switches. To do this we will specify a sequence
$\sigma_{GH}:\quad G=X_{0},X_{1},\ldots,X_{\kappa}=H$
such that $(X_{i},X_{i+1})$ is a $\triangle$-switch for $i=0,\ldots,\kappa-1$. The sequence $\sigma_{GH}$ is called a simulation path for $(G,H)$ of length $\kappa$. Rather than simply prove existence of such a path, we will specify one uniquely for each possible switch $(G,H)$, and show that all our simulation paths have length at most 5. We will do this for degree sequences with minimum degree at least 3, proving the following.
Theorem 1.2.
Let $\boldsymbol{d}$ be a graphical degree sequence with $d_{1}\geq d_{2}\cdots\geq d_{n}\geq 3$ and let $\lambda>0$ be a constant. Then the modified Metropolis $\triangle$-switch chain with parameter $\lambda$ is irreducible on $\mathcal{G}_{n,\boldsymbol{d}}$.
Note that this is not true in general if $d_{n}<3$, see [16]. The above result can be viewed as a reconfiguration result on $\mathcal{G}_{n,\boldsymbol{d}}$, generalising that given in [16]. Reconfiguration problems have been a topic of much recent interest in graph theory. See Nishimura [53] for an introduction. Since Theorem 1.2 depends only on the set of transitions of the Markov chain, and not their probabilities, the value of $\lambda$ is immaterial as long as it is positive.
Our third main result gives a condition under which we can deduce rapid mixing of the modified $\triangle$-switch chain with parameter $\lambda$ from rapid mixing of the switch chain. Recall the definition of $\mu=\mu(\boldsymbol{d})$ from (1).
Theorem 1.3.
Let $\mathcal{D}$ be a family of graphical degree sequences such that for all $\boldsymbol{d}\in\mathcal{D}$ of length $n$ we have $d_{1}\geq d_{2}\geq\cdots\geq d_{n}\geq 3$. Let $\lambda=\lambda(n)\geq 1$ be a function of $n$. Suppose that there exists an $\alpha\in(0,1)$ such that $\lambda\mu(\boldsymbol{d})\leq\log^{\alpha}n$ for every $\boldsymbol{d}\in\mathcal{D}$ of length $n$ and all large enough $n$. If the switch chain is rapidly mixing on $\mathcal{G}_{n,\boldsymbol{d}}$ for all $\boldsymbol{d}\in\mathcal{D}$ then the same is true for the modified $\triangle$-switch chain with parameter $\lambda$.
The switch chain is known to be rapidly mixing for large classes of graphical sequences, in particular regular sequences [14, 30] and all graphical sequences from P-stable families [20]. By Theorem 1.3, the $\triangle$-switch chain is also rapidly mixing for these graphical sequences, whenever $\lambda$ and $\boldsymbol{d}$ satisfy the conditions of the theorem. A special case is stated below. In particular, this corollary implies that if the maximum degree $d_{1}$ is bounded above by a positive constant $C$ for all $\boldsymbol{d}\in\mathcal{D}$ then the $\triangle$-switch chain is rapidly mixing for all $\boldsymbol{d}\in\mathcal{D}$.
Corollary 1.1.
Let $\mathcal{D}$ be a family of graphical sequences such that there exists some constant $\alpha\in(0,1)$ such that $(\log n)^{\alpha/3}\geq d_{1}\geq d_{2}\cdots\geq d_{n}\geq 3$ for all $\boldsymbol{d}\in\mathcal{D}$. Then the modified $\triangle$-switch chain with parameter $\lambda$ is rapidly mixing for any constant $\lambda>1$ and all large enough $n$.
Proof.
Greenhill and Sfragara [30, Theorem 1.1] proved that the switch chain is rapidly mixing for any family $\mathcal{D}$ of degree sequences which satisfy $d_{n}\geq 1$ and $3\leq d_{1}\leq{\textstyle\frac{1}{3}}\sqrt{M}$. These conditions hold under our assumptions when $n$ is large enough. Furthermore,
$\lambda\mu(\boldsymbol{d})\leq\dfrac{\lambda}{6}\,d_{1}^{3}\leq\dfrac{\lambda}{6}\,\log^{\alpha}n\leq\log^{\alpha^{\prime}}n$
for some slightly larger constant $\alpha^{\prime}\in(0,1)$, when $n$ is large enough. Applying Theorem 1.3 completes the proof. ∎
We now describe the structure of the rest of the paper. In Section 1.4 we give some background on the use of canonical paths for bounding the mixing time of Markov chains and describe the two-stage direct canonical path construction method. This machinery is used in Section 1.5 to prove Theorem 1.1. Then in Section 2 we prove the irreducibility result for the $\triangle$-switch chain, Theorem 1.2, and in Section 3 we prove the rapid mixing result for the $\triangle$-switch chain, Theorem 1.3. Finally in Section 4 we will analyse the distribution of the number of triangles under both the uniform distribution on $\mathcal{G}_{n,\boldsymbol{d}}$, and the Gibbs distribution $\pi_{\lambda}$ on $\mathcal{G}_{n,\boldsymbol{d}}$. In particular, the results of Section 4 imply that if $M_{2}\geq M$, the Metropolis switch chain and the modified Metropolis switch chain are polynomial-time indistinguisable under the assumptions of Theorem 1.1, and similarly for the Metropolis $\triangle$-switch chain and the modified Metropolis $\triangle$-switch chain under the assumptions of Theorem 1.3. (The Metropolis $\triangle$-switch chain has transition procedure which is obtained from Figure 4 by replacing $\min\{t(H),\nu\}$ with $t(H)$ and replacing $\min\{t(G),\nu\}$ with $t(G)$.) See Remark 4.
1.4 Background on canonical paths and mixing time
Let $\mathcal{M}$ be an ergodic Markov chain with finite state space $\Omega$, transition matrix $P$ and stationary distribution $\pi$. Let $\mathcal{G}=(\Omega,E(\mathcal{M}))$ be the graph underlying $\mathcal{M}$, so the edge set of $\mathcal{G}$ corresponds to (non-loop) transitions of $\mathcal{M}$.
Given two probability distributions $\sigma$, $\rho$ on $\Omega$, the total variation distance between them, denoted $\mathrm{d}_{\mathrm{TV}}(\sigma,\rho)$, is defined by
$\mathrm{d}_{\mathrm{TV}}(\sigma,\rho)=\tfrac{1}{2}\sum_{x\in\Omega}|\sigma(x)-\rho(x)|=\max_{A\subseteq\Omega}\,(\sigma(A)-\rho(A)).$ (4)
The mixing time $\tau(\varepsilon)$ of $\mathcal{M}$ is given by
$\tau(\varepsilon)=\max_{x\in\Omega}\,\min\,\{t\geq 0\ |\ \mathrm{d}_{\mathrm{TV}}(P^{t}_{x},\pi)\leq\varepsilon\},$
where $P^{t}_{x}$ is the distribution of the random state $X_{t}$ of the Markov chain after $t$ steps with initial state $x$.
Let the eigenvalues of $P$ be
$11=\mu_{0}>\mu_{1}\geq\mu_{2}\cdots\geq\mu_{N-1}>-$ 1
and define $\mu_{\ast}=\max\{\mu_{1},\,|\mu_{N-1}|\}$. Then, see for example [57, Proposition 1], the mixing time of $\mathcal{M}$ satisfies
$\tau(\varepsilon)\leq(1-\mu_{\ast})^{-1}\left(\log(1/\pi^{\ast})+\log(\varepsilon^{-1})\right)$ (5)
where $\pi^{\ast}=\min_{x\in\Omega}\pi(x)$ is the smallest stationary probability. In applications we often focus on the second-largest eigenvalue $\mu_{1}$, if necessary by making the Markov chain lazy (replacing the transition matrix $P$ by $(I+P)/2$ to ensure that all eigenvalues are nonnegative). Alternatively, it follows from a result of Diaconis and Saloff-Coste [19, p. 702] that the smallest eigenvalue of $\mathcal{M}$ satisfies
$(1+\mu_{N-1})^{-1}\leq\max_{x\in\Omega}\frac{1}{2\,P(x,x)}.$ (6)
Jerrum and Sinclair [41] introduced the canonical path method for bounding $\mu_{1}$. For each pair $x,y\in\Omega$, we choose a path $\gamma_{xy}$ from $x$ to $y$ in the graph $\mathcal{G}$ underlying the Markov chain. A canonical choice of paths seeks to produce a small value of the critical parameter $\bar{\rho}(\Gamma)$, the congestion of the chosen set of paths $\Gamma=\{\gamma_{xy}\mid x,y\in\Omega\}$, defined by
$\bar{\rho}(\Gamma)=\max_{e\in E(\mathcal{M})}Q(e)^{-1}\sum_{\begin{subarray}{c}x,y\in\Omega\\ e\in\gamma_{xy}\end{subarray}}\,\pi(x)\pi(y)|\gamma_{xy}|$
where $Q(e)=\pi(u)\,P(u,v)$ when $e=uv\in E(\mathcal{M})$. Then
$(1-\mu_{1})^{-1}\leq\bar{\rho}(\Gamma),$ (7)
so a low congestion can lead to a good bound on the mixing time, using (5). See [57, Theorem 5] or [40].
In [17], the two-stage direct canonical path construction method was described. Here we have two ergodic Markov chains $\mathcal{M}$ and $\mathcal{M}^{\prime}$, with state spaces $\Omega$ and $\Omega^{\prime}$, respectively. We assume that we have a set of canonical paths for $\mathcal{M}^{\prime}$. Furthermore, for each each transition $(Y,Z)$ of $\mathcal{M}^{\prime}$, we specify a (canonical) sequence $\sigma_{YZ}$ of transitions of $\mathcal{M}$ to simulate $(Y,Z)$:
$\sigma_{YZ}:\quad Y=X_{0},X_{1},\ldots,X_{\kappa}=Z$
such that $(X_{i},X_{i+1})$ is a transition of $\mathcal{M}$. We call $\sigma_{YZ}$ a $(\mathcal{M},\mathcal{M}^{\prime})$-simulation path and let $\Sigma=\{\sigma_{YZ}\mid(Y,Z)\in E(\mathcal{M}^{\prime})\}$ be the set of all specified $(\mathcal{M},\mathcal{M}^{\prime})$-simulation paths (one for each transition of $\mathcal{M}^{\prime}$).
For each transition $e\in E(\mathcal{M})$, let
$\Sigma(e)=\{\sigma\in\Sigma\mid e\in\sigma\}$
be the set of all specified simulations paths which use the transition $e$. We define the parameters
$\ell(\Sigma)=\max_{\sigma\in\Sigma}|\sigma|,\quad B(\Sigma)=\max_{e\in E(\mathcal{M})}|\Sigma(e)|$ (8)
representing the maximum length of any simulation path, and the maximum number of simulation paths through a given transition of $\mathcal{M}$, respectively. Finally, the simulation gap between the two chains is defined by
$D(\mathcal{M},\mathcal{M}^{\prime})=\max_{\begin{subarray}{c}uv\in E(\mathcal{M})\\ zw\in E(\mathcal{M}^{\prime})\end{subarray}}\frac{\pi^{\prime}(z)P^{\prime}(z,w)}{\pi(u)P(u,v)}$ (9)
and the stationary ratio is defined by
$R(\mathcal{M},\mathcal{M}^{\prime})=\max_{x,y\in\Omega}\frac{\pi(x)\,\pi(y)}{\pi^{\prime}(x)\pi^{\prime}(y)}=\left(\max_{x\in\Omega}\frac{\pi(x)}{\pi^{\prime}(x)}\right)^{2}.$ (10)
Note that $\pi$ and $\pi^{\prime}$ are both positive on their domains as $\mathcal{M}$, $\mathcal{M}^{\prime}$ are ergodic and reversible.
The following is a slight simplification of [17, Theorem 2.1], tailored to the case that $\Omega=\Omega^{\prime}$. (If $\Omega=\Omega^{\prime}$ then the surjection $h$ mentioned in [17, Theorem 2.1] can be taken to be the identity.) We will also take the opportunity to correct a (fortunately) minor error in [17, Theorem 2.1], where the factor $R(\mathcal{M},\mathcal{M}^{\prime})$ was omitted.
Theorem 1.4.
Let $\mathcal{M}$ and $\mathcal{M}^{\prime}$ be two ergodic Markov chains on a set $\Omega$. Let $\Gamma^{\prime}$ be a set of canonical paths in $\mathcal{M}^{\prime}$ and let $\Sigma$ be a set of $(\mathcal{M},\mathcal{M}^{\prime})$-simulation paths. Then there exists a set of canonical paths in $\mathcal{M}$ whose congestion satisfies
$\bar{\rho}(\Gamma)\leq D(\mathcal{M},\mathcal{M}^{\prime})\,R(\mathcal{M},\mathcal{M}^{\prime})\,\ell(\Sigma)\,B(\Sigma)\,\bar{\rho}(\Gamma^{\prime}).$
Proof.
In the proof of [17, Theorem 2.1] it is shown that
$\bar{\rho}(\Gamma)\leq D(\mathcal{M},\mathcal{M}^{\prime})\,\ell(\Sigma)\,\max_{uv\in E(\mathcal{M})}\,\sum_{\begin{subarray}{c}zw\in E(\mathcal{M}^{\prime})\\ \sigma_{zw}\in\Sigma(uv)\end{subarray}}\frac{1}{\pi^{\prime}(z)P^{\prime}(z,w)}\sum_{\begin{subarray}{c}x,y\in\Omega^{\prime}\\ zw\in\gamma^{\prime}_{xy}\end{subarray}}\pi(x)\pi(y)\,|\gamma^{\prime}_{xy}|$
and that
$\max_{uv\in E(\mathcal{M})}\,\sum_{\begin{subarray}{c}zw\in E(\mathcal{M}^{\prime})\\ \sigma_{zw}\in\Sigma(uv)\end{subarray}}\frac{1}{\pi^{\prime}(z)P^{\prime}(z,w)}\sum_{\begin{subarray}{c}x,y\in\Omega^{\prime}\\ zw\in\gamma^{\prime}_{xy}\end{subarray}}\pi^{\prime}(x)\pi^{\prime}(y)\,|\gamma^{\prime}_{xy}|\leq B(\Sigma)\,\bar{\rho}(\Gamma^{\prime}).$
The result follows as $\pi(x)\pi(y)\leq R(\mathcal{M},\mathcal{M}^{\prime})\,\pi^{\prime}(x)\pi^{\prime}(y)$. ∎
We remark that in the analysis of the flip chain given in [17], the above theorem is applied twice. In one application the missing factor $R$ is exactly 1, and in the other application it can be bounded above by a constant close to 1 (say 2). So the stated bound on the mixing time of the flip chain in [17] is too small by a factor of at most 2, which is of no consequence. However, in the proof of Theorem 1.1 the ratio $R(\mathcal{M},\mathcal{M}^{\prime})$ becomes rather more significant as we compare the uniform distribution to a highly non-uniform distribution.
1.5 Rapid mixing of the modified Metropolis switch chain
We can immediately prove Theorem 1.1 using the two-stage direct canonical path construction method. As well as being of independent interest, this result forms a useful stepping-stone in our proof of Theorem 1.3.
Proof of Theorem 1.1.
Let $\mathcal{M}_{\lambda}$ denote the modified Metropolis switch chain with parameter $\lambda$, as given in Figure 3, and let $\mathcal{M}^{\prime}$ denote the standard switch chain with uniform stationary distribution: that is, $\mathcal{M}^{\prime}$ has transition procedure given by Figure 3 with $\lambda=1$. Suppose that $\mathcal{M}^{\prime}$ is rapidly mixing on all degree sequences in some family $\mathcal{D}$. That is, there exists a polynomial $p(n)$ such that if $\boldsymbol{d}\in\mathcal{D}$ and $\boldsymbol{d}$ has length $n$ then the mixing time of the switch chain on $\mathcal{G}_{n,\boldsymbol{d}}$ is at most $p(n)$.
Using results of Sinclair [57], Guruswami [32, Theorem 4.9] proved that there exists a set $\Gamma^{\prime}$ of canonical paths for $\mathcal{M}^{\prime}$ and a polynomial $q(n)$ such that $\bar{\rho}(\Gamma^{\prime})\leq q(n)$ for all $\boldsymbol{d}\in\mathcal{D}$ with length $n$. Since the switch chain and the modified Metropolis switch chain have the same set of allowable transitions, we may take each simulation path to have length 1. Let $\Sigma$ be the set of these simulations paths. Then $\ell(\Sigma)=1$ and $B(\Sigma)=1$.
The normalising factor $\hat{Z}_{\lambda}(\boldsymbol{d})$ for the stationary distribution of the modified Metropolis switch chain satisfies
$\hat{Z}_{\lambda}(\boldsymbol{d})=\sum_{G\in\mathcal{G}_{n,\boldsymbol{d}}}\lambda^{\min\{t(G),\,\nu\}}\leq\lambda^{\nu}\,|\mathcal{G}_{n,\boldsymbol{d}}|.$ (11)
The switch chain $\mathcal{M}^{\prime}$ has uniform stationary distribution over $\mathcal{G}_{n,\boldsymbol{d}}$ and transition probability $P^{\prime}(G,H)=\frac{1}{3a(\boldsymbol{d})}$ whenever $G\neq H$ differ by a switch. Hence the simulation gap $D(\mathcal{M},\mathcal{M}^{\prime})$ satisfies
$D(\mathcal{M},\mathcal{M}^{\prime})=\max_{uv\in E(\mathcal{M})}\frac{1}{3a(\boldsymbol{d})\,|\mathcal{G}_{n,\boldsymbol{d}}|}\cdot\frac{3a(\boldsymbol{d})\,\hat{Z}_{\lambda}(\boldsymbol{d})}{\min\{\lambda^{\min\{t(u),\nu\}},\lambda^{\min\{t(v),\nu\}}\}}\leq\lambda^{\nu}.$
(Since $\lambda\geq 1$, the maximum over $uv\in E(\mathcal{M})$ is obtained when $t(u)=0$ or $t(v)=0$.) The stationary ratio satisfies
$R(\mathcal{M},\mathcal{M}^{\prime})=\max_{x\in\mathcal{G}_{n,\boldsymbol{d}}}\left(\frac{|\mathcal{G}_{n,\boldsymbol{d}}|\,\lambda^{\min\{t(x),\nu\}}}{\hat{Z}_{\lambda}(\boldsymbol{d})}\right)^{2}\leq\left(\max_{x\in\mathcal{G}_{n,\boldsymbol{d}}}\lambda^{\min\{t(x),\nu\}}\right)^{2}\leq\lambda^{2\nu}.$
Since we take $\nu=\frac{\log n}{\log\log n}$, under the assumptions of the theorem we have
$\lambda^{\nu}\leq\big{(}\log^{\alpha}\big{)}^{\log n/\log\log n}=e^{\alpha\log n}=n^{\alpha}.$ (12)
Hence, by Theorem 1.4, there exists a set $\Gamma$ set of canonical paths for $\mathcal{M}$ (in fact $\Gamma=\Gamma^{\prime}$) with congestion
$\bar{\rho}(\Gamma)\leq n^{3\alpha}\,\bar{\rho}(\Gamma^{\prime}).$
It follows that
$(1-\mu_{1})^{-1}\leq\bar{\rho}(\Gamma)\leq n^{3\alpha}\,q(n).$
Next, the modified Metropolis switch chain has self-loop probability of at least $1/3$ on every state, so (6) implies that
$(1+\mu_{N-1})^{-1}\leq\dfrac{3}{2}.$
Combining the two previous inequalities implies that
$(1-\mu_{*})^{-1}\leq n^{3\alpha}\,q(n).$
To apply (5) we note that under the assumptions of the theorem, the minimum stationary probability for $\mathcal{M}$ satisfies
$\log(1/\pi^{*})=\log\big{(}\hat{Z}_{\lambda}(\boldsymbol{d})\big{)}\leq\log\big{(}\lambda^{\nu}\big{)}+\log\big{(}|\mathcal{G}_{n,\boldsymbol{d}}|\big{)}\leq\log n+M\log M$
using (11). This uses the fact, which can be proved using the configuration model [11], that
$\log|\mathcal{G}_{n,\boldsymbol{d}}|\leq M\log M$ (13)
where $M=M(\boldsymbol{d})=\sum_{j\in[n]}d_{j}$. Hence using (5) we conclude that the modified Metropolis switch chain is rapidly mixing, with mixing time bounded above by
$\tau(\varepsilon)\leq n^{3\alpha}\,q(n)\,\big{(}M\log M+\log n+\log(\varepsilon^{-1})\big{)}.$
This completes the proof of Theorem 1.1. ∎
We can see from this proof why we dampen the impact of the upper tail of the distribution of $t(G)$ by replacing $t(G)$ with $\min\{t(G),\nu\}$ in the transition procedure. The factor $D(\mathcal{M},\mathcal{M}^{\prime})$ involves the ratio $\hat{Z}_{\lambda}(\boldsymbol{d})/|\mathcal{G}_{n,\boldsymbol{d}}|$, which is the expected value of $\lambda^{\min\{t(G),\nu\}}$ under the uniform distribution on $\mathcal{G}_{n,\boldsymbol{d}}$. Without modification, this factor would become the expected value of $\lambda^{t(G)}$ with respect to the uniform distribution on $\mathcal{G}_{n,\boldsymbol{d}}$, and we do not believe that this quantity is polynomially bounded.
2 Irreducibility
In this section we prove that every switch can be simulated by a sequence of at most five $\triangle$-switches.
Theorem 2.1.
Let $(G,H)$ be a switch, where $G,H\in\mathcal{G}_{n,\boldsymbol{d}}$. There is a $\triangle$-switch simulation path of length at most 5 which simulates the switch $(G,H)$.
We will present the constructions diagrammatically, with the following conventions: solid lines represent known edges of $G$, dotted lines represent known non-edges, and dashed lines represent non-edges involved in the current $\triangle$-switch.
Suppose that a $\triangle$-switch $(G,H)$ replaces the edges $a_{1}a_{2}$,  $a_{3}a_{4}$ with the edges $a_{1}a_{3}$,  $a_{2}a_{4}$. We classify switches into three types, depending on whether one or both of the diagonals $a_{1}a_{4}$, $a_{2}a_{3}$ is present or absent, as shown in Fig. 5. Note that the diagonals have the form $a_{j}a_{5-j}$ for $j=1,2$.
$a_{1}$$a_{2}$$a_{3}$$a_{4}$ype (a)$a_{1}$$a_{2}$$a_{3}$$a_{4}$ype (b)$a_{1}$$a_{2}$$a_{3}$$a_{4}$ype (c)
Fig. 5: The three types of switch, depending on how many diagonals are present
Let $D=\{a_{1},a_{2},a_{3},a_{4}\}$ be the vertices of the switch, and $A_{i}=\mbox{N}(a_{i})\setminus D$ be the neighbours of $a_{i}$ outside of $D$, for $i=1,\ldots,4$. Note that $|A_{i}|=\deg(a_{i})-\deg_{D}(a_{i})$, so $|A_{i}|\geq d_{n}-1$ or $|A_{i}|\geq d_{n}-2$. Thus, since $d_{n}\geq 3$, we have $A_{i}\neq\emptyset$ for all $i\in[4]$ in any switch.
Lemma 2.1.
Suppose that $(G,H)$ is a switch which removes the edges $a_{1}a_{2}$, $a_{3}a_{4}$ and replaces them with the edges $a_{1}a_{3}$, $a_{2}a_{4}$. Further suppose that at least one of the following holds:
  • (i)
    The switch is Type (b) or Type (c). That is, at least one diagonal is present.
  • (ii)
    The switch is Type (a), and some pair of distinct elements of $D=\{a_{1},a_{2},a_{3},a_{4}\}$ have a common neighbour in $[n]\setminus D$.
  • (iii)
    There is a triangle on vertices $a_{j},u,w$ for some vertices $u,w$ which satisfy $u\not\in A_{5-j}$ and $\{u,w\}\cap D=\emptyset$, where $j\in[4]$.
Then there is a $\triangle$-switch simulation path of length at most 4 which simulates the switch $(G,H)$.
Proof.
For future reference, we will number the cases which arise when defining the simulation paths. This will be needed later in the proof of rapid mixing. We treat these cases in order, so we only move to (II) if (I) does not hold, and so on.
  • (I)
    If the switch $(G,H)$ is itself a $\triangle$-switch then the simulation path is $G,H$ which has length 1. This occurs if and only if at least one of $A_{1}\cap A_{2}$, $A_{1}\cap A_{3}$, $A_{3}\cap A_{4}$, $A_{2}\cap A_{4}$ is nonempty.
  • (II)
    Now suppose that the switch has Type (a) (both diagonals absent) and $(A_{1}\cap A_{4})\cup(A_{2}\cap A_{3})$ is non-empty. After relabelling if necessary, assume that $A_{1}\cap A_{4}\neq\emptyset$. Then there is a simulation path of length 2, shown below.
[…]
  • (III)
    Next, suppose that the switch has Type (b) (both diagonals present) and $(A_{1}\cap A_{4})\cup(A_{2}\cap A_{3})$ is non-empty. After relabelling if necessary, $A_{1}\cap A_{4}\neq\emptyset$. Again, there is a simulation path of length 2, as shown below.
[…]
  • (IV)
    Note that if none of (I), (II) or (III) hold then $A_{i}\cap A_{j}=\emptyset$ for all distinct $i,j\in[4]$. Next we suppose that, after relabelling if necessary,
    • at least one of the diagonal edges $a_{j}a_{5-j}$ is present, and the symmetric difference $A_{i}\oplus A_{5-i}$ is nonempty, where $\{i,j\}=\{1,2\}$, or
    • there is a triangle on vertices $a_{j},u,w$ for two distinct vertices $u,w\in[n]\setminus D$. (Here we do not care how many diagonal edges are present.)
    In the first subcase, choose $(i,u)$ to be lexicographically-least. Then after relabelling if necessary, we can assume that $u$ is the lexicographically-least element of $A_{1}\setminus A_{4}$. In the second subcase, choose $(j,u,w)$ to be the lexicographically-least triple satisfying the requirements. Then after relabelling if necessary, we can assume that $j=1$. In both cases, we have a simulation path of length 2, consisting of first replacing edges $a_{1}u,a_{3}a_{4}$ by $a_{1}a_{3},ua_{4}$ and then replacing edges $ua_{4},a_{1}a_{3}$ by edges $a_{1}u,a_{2}a_{4}$. Under the given conditions, each of these operations is a $\triangle$-switch, as required. We illustrate the simulation path below for the first subcase when exactly one diagonal $a_{2}a_{3}$ is present (noting that it this still gives a sequence of two $\triangle$-switches if $a_{1}a_{4}$ is also present), and in the second subcase when neither diagonal is present (noting that this still gives a sequence of two $\triangle$-switches if one or both diagonals are present).
[…]
  • (V)
    Next, suppose that the switch has Type (c), so exactly one diagonal $a_{j}a_{5-j}$ is present, and that $A_{i}=A_{5-i}$, and that $G[A_{i}]$ has at least one edge, where $\{i,j\}=\{1,2\}$. By relabelling if necessary, we may assume that diagonal $a_{1}a_{4}$ is present, and for ease of notation let $A=A_{2}=A_{3}$. Let $(u,v)$ be the lexicographically-least pair of vertices in $A$ such that $uv$ is an edge of $G$. Then $a_{2},u,v$ and $a_{3},u,v$ both determine triangles, while $u\not\in A_{1}$ and $v\not\in A_{4}$ as we are not in (I). Then there is a simulation path of length 4 as shown below.
[…]
  • (VI)
    Finally, suppose that the switch has Type (c), so exactly one diagonal $a_{j}a_{5-j}$ is present, and that $A_{i}=A_{5-i}$, and that $G[A_{i}]$ has no edges, where $\{i,j\}=\{1,2\}$, Again, by relabelling if necessary, we assume that diagonal $a_{1}a_{4}$ is present, and write $A=A_{2}=A_{3}$. Note that $|A|\geq 2$ as $d_{D}(a_{2})=d_{D}(a_{3})=1$. Let $(u,v)$ be the lexicographically-least pair of distinct elements in $A$, and let $w$ be the smallest-labelled vertex in $\mbox{N}(v)\setminus D$, which exists as the minimum degree is at least 3. Then $w\not\in A=A_{3}$ as $v\in A$ and $vw\in E$. We can construct the following simulation path of length 4.
[…]
To complete the proof, observe that (I), (III) and (IV) together cover all Type (b) switches, while Type (c) switches are covered by combining (I), (IV), (V) and (VI). This proves the lemma under assumption (i). The proof for assumption (ii) is covered by (I) and (II) together, while (IV) completes the proof for assumption (iii). ∎
For the remainder of the proof we assume that the assumptions of Lemma 1 do not hold. Hence the switch has Type (a), that is, no diagonals are present. Furthermore, no two vertices in $D$ have a common neighbour outside $D$, and there is no triangle of the form $a_{j}uw$ with $\{u,w\}\cap D=\emptyset$, where $j\in[4]$.
In this situation it is possible that the switch edges $a_{1}a_{2}$, $a_{3}a_{4}$ lie in different components of $G$, so that $H$ has one fewer components that $G$. Only a $\triangle^{-}$-switch can reduce the number of components, and so our strategy is to “plant” a triangle of the form $a_{j},u,w$ with $u,w\not\in D$ distinct. By assumption, $ua_{5-j}$ is not an edge as $A_{j}\cap A_{5-j}=\emptyset$. Hence, after planting the triangle, we can simulate the switch using the simulation path described in part (IV) of Lemma 2.1, then “unplant” the planted triangle. There is one exceptional situation where we will not plant a triangle, but instead will simulate the switch $(G,H)$ with a simulation path of length three. (This is case (IXa) of Lemma 2.3 below.)
We will make use of the following lemma to plant a triangle on some element of $D$. This lemma will be applied to graphs in $\mathcal{G}_{n,\boldsymbol{d}}$, or to graphs in $\mathcal{G}_{n,\boldsymbol{d}}$ with one edge removed.
Lemma 2.2.
Let $G$ be a graph on the vertex set $[n]$. Suppose $v\in[n]$ has a set of three distinct neighbours $R=\{r_{1},r_{2},r_{3}\}$, with $\deg(r_{i})\geq 2$ for $i=1,2,3$. Assume that there is no triangle in $G$ which contains $v$ and an element of $R$.
  1. (i)
    Then there is a $\triangle^{+}$-switch $(G,\widetilde{G})$ such that $\widetilde{G}$ contains a triangle $T$ with $v\in T$ and $R\cap T\neq\emptyset$.
  2. (ii)
    Furthermore, we can insist that $r_{1}\in T$ unless for $j=2,3$ there is a 5-cycle which contains the path $r_{1}vr_{j}$, but no 4-cycle in $G$ contains this path.
Proof.
Let $\ell$ be the length of the shortest cycle in $G$ which contains a path of the form $r_{i}vr_{j}$, where $r_{i},r_{j}$ are distinct elements of $R$. If no such cycle exists, set $\ell=\infty$. By assumption, $\ell\geq 4$. First suppose that $\ell=4$. After relabelling the vertices in $R$ if necessary, assume that the cycle contains the path $r_{1}vr_{2}$ together with some vertex $w$. (For definiteness, let $w$ be the least-labelled common neighbour of $r_{1}$ and $r_{2}$.) Note that $vw$ and $r_{2}r_{3}$ are not edges of $G$, by assumption. Then we can plant a triangle on $v,r_{1},w$ as follows:
[…]
Next suppose that $\ell=5$. After relabelling elements of $R$ if necessary, we can assume that the 5-cycle is $vr_{1}w_{1}w_{2}r_{2}v$, where $w_{1},w_{2}\not\in\{v\}\cup R$ are distinct. For definiteness, choose $(w_{1},w_{2})$ to be lexicographically least among all possibilities. Note that $vw_{1}$ is a non-edge in $G$ by assumption, and $w_{2}r_{3}$ is a non-edge in $G$ as $\ell\neq 4$. We can “plant” a triangle as follows:
[…]
Now suppose that $\ell\geq 6$. After relabelling vertices in $R$ if necessary, suppose that we have a path $w_{1}r_{1}vr_{2}w_{2}$ in $G$, where $w_{1},w_{2}\not\in\{v\}\cup R$ are distinct. (For definiteness, choose $(w_{1},w_{2})$ to be lexicographically least among all options.) Note that $r_{1}r_{2}$ is a non-edge in $G$ by assumption, and $w_{1}w_{2}$ is a non-edge in $G$ as $\ell\neq 5$. Then we can plant a triangle as follows:
[…]
In all cases, we have (uniquely) identified a $\triangle^{+}$-switch which produces a graph $\widetilde{G}$ with the same degree sequence as $G$ which contains a triangle satisfying the stated conditions. This proves (i).
To prove the second statement, adapt the above proof by letting $\ell$ be the length of the shortest cycle which contains the path $r_{1}vr_{j}$ for some $j\in\{2,3\}$. If $\ell\neq 5$ then the $\triangle$-switch described above creates a triangle $T$ with $r_{1}\in T$, as required. If $\ell=5$ then the $\triangle$-switch described above works if and only if $w_{2}r_{3}$ is not an edge, assuming (after swapping the labels of $r_{2}$ and $r_{3}$ if necessary) that the 5-cycle $vr_{1}w_{1}w_{2}r_{2}v$ is present. This completes the proof of (ii). ∎
To describe how to remove a planted triangle, we need some terminology. Let $(G,\widetilde{G})$ be a $\triangle$-switch, where $\widetilde{G}$ is obtained from $G$ by deleting the edges $e_{1},e_{2}$ and inserting the edges $e_{3},e_{4}$. If $\widetilde{H}$ is any graph which contains the edges $e_{3},e_{4}$ and does not contain $e_{1},e_{2}$, then the inverse of the $\triangle$-switch $(G,\widetilde{G})$, applied to $\widetilde{H}$, is the $\triangle$-switch $(\widetilde{H},H)$ where $H$ is obtained from $\widetilde{H}$ by deleting the edges $e_{3},e_{4}$ and inserting the edges $e_{1},e_{2}$. This matches the usual notion of an inverse operation when $\widetilde{H}=\widetilde{G}$, since then $H=G$, but here we need to apply inverse $\triangle$-switches when $\widetilde{H}\neq\widetilde{G}$.
For us, $(G,\widetilde{G})$ will be a $\triangle^{+}$-switch which plants a triangle. Then we will perform two or three $\triangle$-switches from $\widetilde{G}$ to reach a graph $\widetilde{H}$ which contains $e_{3},e_{4}$ but not $e_{1},e_{2}$. Finally we will remove the planted triangle by applying the inverse of $(G,\widetilde{G})$ to $\widetilde{H}$. There is one exception, which is case (IXa) below, where we do not plant a triangle but instead simulate the required switch with a simulation path of length three.
Lemma 2.3.
Suppose that $(G,H)$ is a switch which removes the edges $a_{1}a_{2}$, $a_{3}a_{4}$ and replaces them with the edges $a_{1}a_{3}$, $a_{2}a_{4}$. Further suppose that the following conditions all hold:
  • The switch has Type (a) (no diagonals present),
  • no pair of vertices in $D=\{a_{1},a_{2},a_{3},a_{4}\}$ have a common neighbour, and
  • there is no triangle on vertices $a_{j},u,w$ with $\{u,w\}\cap D=\emptyset$, where $j\in[4]$.
Then there is a $\triangle$-switch simulation path of length at most five which simulates the switch $(G,H)$.
Proof.
We will number the cases which arise when defining the simulation paths. This will be needed later in the proof of rapid mixing. Since Lemma 2.1 has cases (I)–(VI), here our numbering starts from (VII). We treat these cases in order, so we only move to (VIII) if (VII) does not hold, and so on.
  • (VII)
    First suppose that after at least one of $a_{1},a_{2},a_{3},a_{4}$ has degree at least 4. Choose the least-labelled element of $D$ which has degree at least 4: after relabelling the elements of $D$ if necessary, we can assume that $a_{1}$ has degree at least 4. We apply Lemma 2.2(i) with $v=a_{1}$, and $R$ a set of three neighbours of $a_{1}$ other than $a_{2}$. This identifies a $\triangle^{+}$-switch $(G,\widetilde{G})$ where $\widetilde{G}$ contains a triangle of the form $a_{1},u,w$. From $\widetilde{G}$ we can perform the two $\triangle$-switches described in Lemma 2.1, part (IV), to delete the edges $a_{1}a_{2},a_{3}a_{4}$ and replace them with $a_{1}a_{3}$, $a_{1}a_{4}$ using two $\triangle$-switches. From the resulting graph $\widetilde{H}$ we apply the inverse of the $\triangle^{+}$-switch $(G,\widetilde{G})$. This “unplants” the planted triangle and takes us to $H$. This gives a simulation path of length 4.
  • (VIII)
    Next, assume that $\deg(a_{i})=3$ for all $i\in[4]$ and that there is a path of length at most 2 in $G\setminus D$ between some element of $A_{j}$ and some element of $A_{2+j}$, for some $j\in\{1,2\}$. Take a shortest such path, and by relabelling if necessary assume that $j=1$. Let $A_{1}=\{u_{1},u_{2}\}$ and $A_{3}=\{v_{1},v_{2}\}$ and by relabelling these vertices if necessary, assume that the path has endvertices $u_{1}$, $v_{1}$. Note that $u_{2}\not\in A_{4}$ and $v_{1}\not\in A_{2}$, since no two elements of $D$ have a common neighbour (by assumption).
    (VIIIa): If the path has length 1 then the edge $u_{1}v_{1}$ is present and we can plant a triangle as follows.
[…]
    We have a triangle $a_{1},u_{1},v_{2}$ which allows us to perform the two $\triangle$-switches as described in Lemma 2.1, part (IV), to simulate the required switch. From the resulting graph, we apply the inverse of the triangle-planting $\triangle$-switch, to “unplant” the triangle and produce $H$. This gives a simulation path of length 4.
    (VIIIb): If the path has length 2 then (after possibly relabelling as described above), $u_{1}$ and $v_{1}$ have a common neighbour $w$, where we choose $w$ to be the least-labelled common neighbour of $u_{1}$, $v_{1}$. Note that $a_{3}w\notin E$ because $\deg(a_{3})=3$, and $u_{2}v_{1}\notin E$ since otherwise it would be an edge between the neighbourhoods of $a_{1},a_{3}$. Then we can plant a triangle as follows:
[…]
    Now we have a triangle $a_{1},u_{1},w$ which we can use to perform the two $\triangle^{+}$-switches as specified in Lemma 2.1, case (IV). From the resulting graph, we apply the inverse of the triangle-planting $\triangle$-switch, to “unplant” this triangle and produce $H$. This gives a simulation path of length 4.
  • (IX)
    Here we assume that $\deg(a_{i})=3$ for all $i\in[4]$ and that there is no path of length at most 2 between an element of $A_{j}$ and an element of $A_{2+j}$, for $j=1,2$. Let $A_{1}=\{u_{1},u_{2}\}$.
    (IXa) First suppose that there is an edge between two neighbours of $u_{j}$, and by relabelling if necessary we can assume that $j=1$. Then $w_{1}w_{2}\in E$ for some $w_{1},w_{2}\in\mbox{N}(u_{1})$ which are distinct from $a_{1}$. (For definiteness, let $(w_{1},w_{2})$ be the lexicographically-least such pair.) In this subcase, we do not plant a triangle: instead we use the following simulation path of length 3, where $v$ is the least-labelled element of $A_{3}$. Note that $u_{1}a_{3}\not\in E$ since $A_{1}\cap A_{3}=\emptyset$, and $vw_{2}\not\in E$ since $u_{1}$ and $v$ are at distance at least 3 in $G$.
[…]
    (IXb) Now suppose that there is no edge between any two neighbours of $u_{j}$ for $j\in\{1,2\}$. Let $w_{1},w_{2}$ be the two least-labelled neighbours of $u_{1}$ distinct from $a_{1}$. We want to apply Lemma 2.2(ii) to the graph $G^{(-)}:=G-\{a_{1}a_{2}\}$ with the edge $a_{1}a_{2}$ deleted, setting $v=u_{1}$ and $R=\{a_{1},w_{1},w_{2}\}$ with $r_{1}=a_{1}$. Suppose that Lemma 2.2(ii) succeeds in identifying a $\triangle^{+}$-switch $(G^{(-)},\widetilde{G}^{(-)})$ such that $\widetilde{G}^{(-)}$ contains a triangle on the vertices $a_{1},u_{1},z$ for some vertex $z\not\in D$. Let $\widetilde{G}$ denote the graph on $\mathcal{G}_{n,\boldsymbol{d}}$ obtained by inserting the edge $a_{1}a_{2}$ in $\widetilde{G}^{(-)}$. Then $(G,\widetilde{G})$ is a $\triangle$-switch which can be applied to $G$. From $\widetilde{G}$, we can perform the two $\triangle$-switches described in Lemma 2.1 case (IV), to exchange the switch edges. Then from the resulting graph, apply the inverse of the $\triangle$-switch $(G,\widetilde{G})$ to “unplant” the triangle and produce the graph $H$. This gives a simulation path of length 4.
    (IXc) Finally we assume that there is no edge between any two neighbours of $u_{j}$, for $j\in\{1,2\}$, but that Lemma 2.2(ii) did not succeed when applied to the graph $G^{(-)}:=G-\{a_{1}a_{2}\}$ with $v=u_{1}$, $R=\{a_{1},w_{1},w_{2}\}$ and $r_{1}=a_{1}$. (As above, we assume that $w_{1},w_{2}$ are the two least-labelled neighbours of $u_{1}$ distinct from $a_{1}$.) In particular, this implies that $G^{(-)}$ has a 5-cycle $C$ which contains the path $a_{1}u_{1}w_{1}$, but $G^{(-)}$ does not have any 4-cycle which contains the path $a_{1}u_{1}w_{j}$ for $j\in\{1,2\}$. The same properties are true for $G$. Since $a_{1}$ has degree 2 in $G^{(-)}$, the 5-cycle $C$ must be $a_{1}u_{1}w_{1}zu_{2}$ for some vertex $z$. Note that $u_{1}z$ cannot be an edge of $G$, or we are in case (IXa), and $w_{2}u_{2}$ cannot be and edge of $G$ as this would give rise to a 4-cycle $a_{1}u_{1}w_{2}u_{2}$ in $G^{(-)}$ containing the path $a_{1}u_{1}w_{2}$. Hence we can perform the following $\triangle$-switch in $G$ to plant a triangle on $u_{1},w_{1},z$.
[…]
    Then we have the situation in (IXa), where $u_{2}$ and its neighbours are temporarily unimportant. The three $\triangle$-switches given in case (IXa) perform the required switch, and from the resulting graph we apply the inverse of the $\triangle$-switch which planted $u_{1}w_{1}w_{2}$, reaching the graph $H$. This gives a simulation path of length five.∎
Now Theorem 2.1 follows immediately by combining Lemma 2.1 and Lemma 2.3. Finally, Theorem 1.2 is obtained from Theorem 2.1 using the fact that the switch chain is irreducible on $\mathcal{G}_{n,\boldsymbol{d}}$ for all graphical degree sequences $\boldsymbol{d}$ (see [23]).
3 Rapid mixing of the modified Metropolis $\triangle$-switch chain
In this section we prove Theorem 1.3 using the two-stage direct canonical path construction method, restated above as Theorem 1.4. Recall the definitions from Section 1.4. Let $\mathcal{M}$ denote the modified Metropolis $\triangle$-switch chain (see Figure 4) and let $\mathcal{M}^{\prime}$ denote the modified Metropolis switch chain (see Figure 3).
In Section 2 we constructed a set
$\Sigma=\{\sigma_{GH}\mid(G,H)\in E(\mathcal{M}^{\prime})\}$
of $(\mathcal{M},\mathcal{M}^{\prime})$-simulation paths. Recall the parameters defined in (8). First we prove an upper bound on the parameter $B(\Sigma)$, which measures the maximum number of simulation paths containing any given $\triangle$-switch.
Lemma 3.1.
Let $\boldsymbol{d}$ be a graphical degree sequence with $d_{1}\geq d_{2}\geq\cdots\geq d_{n}\geq 3$ and let $\Sigma$ be the set of $(\mathcal{M},\mathcal{M}^{\prime})$-simulation paths defined in Section 2. Then
$B(\Sigma)\leq 20d_{1}^{2}\,(2M+d_{1}^{2})$
where $M=M(\boldsymbol{d})=\sum_{j\in[n]}d_{j}$.
Proof.
Let $e=(X,Y)$ be a transition of the $\triangle$-switch chain. We want an upper bound on the number of simulation paths $\sigma_{GH}$ which contain $e$, where $(G,H)$ is a transition of the switch chain. To do this, we will work through the cases (I)–(IX) from Lemmas 2.1 and 2.3. We will not try to optimise the bounds we determine. We are only interested here in correctness and, to a lesser extent, simplicity.
Let $D=\{a_{1},a_{2},a_{3},a_{4}\}$ denote the vertices involved in the switch $(G,H)$ that we need to identify.
  • (I)
    In Case (I) we have $(G,H)=(X,Y)$ which happens in exactly 1 way.
  • (II)
    In Case (II), the simulation path has length 2 and only involves the vertices in $D$. So the simulation path is $G,G_{1},H$ and in $G_{1}$, the edges $a_{1}a_{4}a_{2}a_{3}$ are present and the switch edges $a_{1}a_{2}$, $a_{1}a_{3}$, $a_{2}a_{4}$, $a_{3}a_{4}$ are all absent. Therefore, if $(X,Y)=(G,G_{1})$ then we know $G=X$ and $H$ is obtained from $G_{1}$ by replacing edges $a_{1}a_{4}$, $a_{2}a_{3}$ with $a_{1}a_{3}$, $a_{2}a_{4}$ (the only perfect matching of $D$ not seen in either $G$ or $G_{1}$). This can be done in exactly one way. By a similar argument, if $(X,Y)=(G_{1},H)$ then there is exactly one way to reconstruct $G$. So overall, there are 2 possibilities for $(G,H)$ if $(X,Y)$ is involved in a simulation path in Case (II).
  • (III)
    There are exactly 2 possibilities arising from Case (III), using a very similar argument to that used for Case (II).
  • (IV)
    The simulation path defined in Case (IV) has length 2, so it is $G,G_{1},H$ for some $G_{1}$. Each $\triangle$-switch in the path involves 3 vertices in $D$ and one vertex not in $D$. (Note that in the second subcase, the vertex $w$ of the triangle on $a_{1},u,w$ is not involved in either of the $\triangle$-switches in the simulation path.) There are 4 ways to label the vertices involved in a given $\triangle$-switch, and in each case the remaining vertex of $D$ is a neighbour of one of the four vertices involved in the $\triangle$-switch. Hence there are at most $4d_{1}$ ways to label all vertices in $D$ or involved in the $\triangle$-switch, and using these vertices we can uniquely construct $H$ from $(G,G_{1})$, or uniquely construct $G$ from $(G_{1},H)$. Hence there are at most $8d_{1}$ possibilities for $(G,H)$ if $(X,Y)$ is involved in a simulation path in Case (IV).
  • (V)
    The simulation path in Case (V) has length 4, so it is $G,G_{1},G_{2},G_{3},H$ for some graphs $G_{1},G_{2},G_{3}$. If $(X,Y)=(G,G_{1})$ then one of the edges switched out is the diagonal $a_{1}a_{4}$, the the other edge switched out is $uv$ where $u,v\not\in D$. Then $a_{2}$ is a common neighbour of of $a_{1},u,v$ in $G$, and $a_{3}$ is a common neighbour of $a_{4},u,v$ in $G$. So there are 2 choices for which of the removed edges is $a_{1}a_{4}$ and at most $d_{1}^{2}$ ways to choose the unordered pair $\{a_{2},a_{3}\}$. From $G=X$ and this information we can recover $H$. (Note that by symmetry we do not need to multiply by a factor of 2 to decide which vertex of the edge $a_{1}a_{4}$ is $a_{1}$. The same argument works for the final $\triangle$-switch in the simulation path.
    Next, if $(X,Y)=(G_{1},G_{2})$ then the $\triangle$-switch involves the vertices $a_{1},a_{2},a_{3},a_{4}$, removing the edges $a_{1}a_{2}$, $a_{3}a_{4}$ which are present in $G$ and inserting the diagonals $a_{1}a_{4}$, $a_{2}a_{3}$. Hence we know that $H[D]$ is given by the only perfect matching of $D$ not seen in either $X$ or $Y$. To recover $G$ from $X$ we must know the vertices $u,v$ which are both common neighbours of $a_{2}$ and $a_{3}$, so there are at most $d_{1}^{2}$ choices for $u,v$. The same argument holds if $(X,Y)=(G_{2},G_{3})$. Overall, there are at most $6d_{1}^{2}$ ways that $(X,Y)$ could be involved in a simulation path in Case (V).
  • (VI)
    The simulation path in Case (VI) has length 4. Each $\triangle$-switch in this simulation path involves four of the vertices $a_{1},a_{2},a_{3},a_{4},u,v,w$. There are 4 ways to label the vertices involved in each $\triangle$-switch in this path, and then the remaining 3 vertices can be chosen in at most $d_{1}^{3}$ ways, since the subgraph induced by $\{a_{1},a_{2},a_{3},a_{4},u,v,w\}$ is connected at each step. Once the identities of these vertices are known, from any $\triangle$-switch on the simulation path, we can construct $(G,H)$ uniquely. It follows that there are at most $16d_{1}^{3}$ ways that $(X,Y)$ could be involved in a simulation path in Case (VI).
  • (VII)
    The simulation path in Case (VII) has length 4, consisting of a triangle-planting $\triangle$-switch, then the simulation path from Case (IV) in the resulting graph, then the “unplanting” $\triangle$-switch. Lemma 2.2 is used to choose the triangle-planting $\triangle$-switch, with $v=a_{1}$ and $R\subseteq\mbox{N}(a_{1})\setminus D$. If $(X,Y)$ is the triangle-planting $\triangle$-switch $(G,\widetilde{G})$ then $X=G$. If (in the application of Lemma 2.2) $\ell=4$ then the triangle-planting $\triangle$-switch involves one vertex from $D$ and three vertices outside $D$. There are 4 ways to label the vertices involved in the $\triangle$-switch, which identifies $a_{1}$ (say). Then there at most $d_{1}$ ways to choose $a_{2}$ from the neighbours of $a_{1}$, and at most $M$ ways to choose the vertices $(a_{3},a_{4})$. Using these vertices we can construct $H$. So there are at most $4d_{1}M$ ways that $(X,Y)$ can be the triangle-planting $\triangle$-switch when $\ell=4$, and the same bound holds when $\ell=5$.
    However, if $\ell\geq 6$ then the triangle-planting $\triangle$-switch does not involve $a_{1}$, but involves two neighbours $u_{1},u_{2}$ of $a_{1}$ (the neighbours form a triangle with $a_{1}$ after the $\triangle$-switch). Here there are 2 choices for the edge $u_{1}u_{2}$ which is inserted by the $\triangle$-switch, then at most $d_{1}^{2}$ choices for $(a_{1},a_{2})$ and at most $M$ choices for $(a_{3},a_{4})$. Once these vertices are known, we can construct $H$ from $G$. There are at most $2d_{1}^{2}M$ ways that $(X,Y)$ can be the triangle-planting switch when $\ell\geq 6$. Hence, overall there are at most $8d_{1}M+2d_{1}^{2}M$ ways that $(X,Y)$ could be the triangle-planting $\triangle$-switch in Case (VII). The same bound holds for the “unplanting” $\triangle$-switch at the end of the simulation path, and we must also add the $8d_{1}$ choices that arise if $(X,Y)$ is used as the second or third $\triangle$-switch in the simulation path. (We multiply $8d_{1}$ by three as there are three possibilities for $\ell$, as described above.) As a grand total, there are at most
    $4d_{1}^{2}M+16d_{1}M+24d_{1}$
    possiblities for $(G,H)$ if $(X,Y)$ is part of a simulation path in Case (VII).
  • (VIII)
    In subcase (VIIIa), the triangle-planting $\triangle$-switch involves two vertices from $D$, say $a_{1},a_{3}$, and two vertices outside $D$. There are 4 ways to label the vertices involved in the $\triangle$-switch and then at most $d_{1}^{2}$ ways to choose $(a_{2},a_{4})$. So if $(X,Y)$ is the triangle-planting $\triangle$-switch in subcase (VIIIa) then there are at most $4d_{1}^{2}$ options for $(G,H)$, as $G=X$ and $H$ can be constructed from $G$ using these known vertices. The same bound holds for the unplanting $\triangle$-switch in subcase (VIIIa), and again we have at most $8d_{1}$ options from the second or third $\triangle$-switch in the simulation path, giving at most $8d_{1}^{2}+8d_{1}$ choices for $(G,H)$ from subcase (VIIIa).
    In subcase (VIIIb), the triangle-planting $\triangle$-switch involves exactly one vertex from $D$, say $a_{1}$, and three vertices outside $D$. One of the vertices involved in the triangle-planting $\triangle$-switch is a neighbour of $a_{3}$, so there are at most $4d_{1}^{3}$ ways to identify all vertices in the triangle-planting $\triangle$-switch and all vertices in $D$. The same bound holds for the unplanting $\triangle$-switch, and again we must include the bound $8d_{1}$ to cover the second and third $\triangle$-switches in the simulation path. Hence there are at most $8d_{1}^{3}+8d_{1}$ choices for $(G,H)$ if $(X,Y)$ is part of a simulation path in subcase (VIIIb).
    Overall, there are at most
    $8d_{1}^{3}+8d_{1}^{2}+16d_{1}$
    choices for $(G,H)$ if $(X,Y)$ is part of a simulation path in case (VIII).
  • (IX)
    In subcase (IXa), the simulation path has length 3, so it is $G,G_{1},G_{2},H$ for some graphs $G_{1},G_{2}$. (Here we do not plant a triangle.) The first $\triangle$-switch involves one vertex of $D$, say $a_{3}$, and a neighbour of $a_{1}$. So there are at most $4d_{1}^{3}$ ways to identify all vertices in the $\triangle$-switch as well as $(a_{1},a_{2},a_{4})$. Here $G=X$ and we can construct $H$ using the vertices already identified. The second $\triangle$-switch $(G_{1},G_{2})$ involves the four vertices of $D$, with $u_{1}$ a common neighbour of $a_{1}$ and $a_{3}$ (say) and where $u_{1}w_{1}w_{2}v$ is a path in $G_{1}$. Hence there are at most $2d_{1}^{4}$ ways to identify each of the vertices of $D\cup\{u_{1},w_{1},w_{2}\}$. Using this we can construct $G$ from $G_{1}$ and then construct $H$. Finally, the third switch involves exactly one vertex from $D$, say $a_{3}$, but now it is possible that $a_{2}$ and $a_{4}$ belong to a different component than $a_{3}$ in both $G_{2}$ and $H$. Since $H=Y$, there are at most $4d_{1}M$ ways to identify the vertices of $D$ and then construct $G$ from $H$. Overall there are most $4d_{1}M+2d_{1}^{4}+4d_{1}^{3}$ choices for $(G,H)$ if $(X,Y)$ is part of a simulation path in subcase (IXa).
    In subcase (IXb), we call Lemma 2.2 to identify a triangle-planting switch $(G,\widetilde{G})$. Then from $\widetilde{G}$ we perform two $\triangle$-switches as described in Case (IV), then we “unplant” the triangle by performing the inverse $\triangle$-switch in the resulting graph. There are several possibilities which we must analyse:
    • If Lemma 2.2 uses a 4-cycle which includes the path $a_{1}u_{1}w_{j}$ for $j\in\{1,2\}$, then the triangle-planting $\triangle$-switch does not involve any vertex of $D$, but involves a neighbour of $a_{1}$, say. Hence there are at most $4d_{1}^{2}M$ ways to identify the vertices of $D$. Since $G=X$, this allows us to construct $H$.
    • If $a_{1}u_{1}w_{j}$ is not part of a 4-cycle or 5-cycle ($j=1,2$) then the triangle-planting $\triangle$-switch involves $a_{1}$, and hence there are at most $4d_{1}M$ ways to identify the vertices of $D$ and hence construct $H$ from $G=X$.
    • If $a_{1}u_{1}w_{j}$ is in a 5-cycle but is not in any 4-cycle, then the triangle-planting $\triangle$-switch does not involve any element of $D$ but it involves neighbours of $a_{1}$, say. So there are at most $4d_{1}^{2}M$ ways to identify the vertices of the $\triangle$-switch and of $D$, which allows us to construct $H$ uniquely from $G=X$.
    In each case, the same bounds hold for the final “unplanting” $\triangle$-switch, and there are at most $8d_{1}$ choices for $(G,H)$ which arise if $(X,Y)$ is the second or third $\triangle$-switch in the simulation path. (We multiply $8d_{1}$ by 3 to account for the three possible situations for $\ell$.) In total, this gives at most
    $2(8d_{1}^{2}M+4d_{1}M)+24d_{1}=16d_{1}^{2}M+8d_{1}M+24d_{1}$
    choices for $(G,H)$ when $(X,Y)$ is part of a simulation path in subcase (IXb).
    In subcase (IXc), we require a $\triangle$-switch to plant a triangle, say $(G,\widetilde{G})$. The $\triangle$-switch involves no vertex from $D$ but involves a neighbour of one vertex from $D$, say $a_{1}$. Here $G=X$ and there are at most $4d_{1}^{2}M$ ways to identify the elements of the $\triangle$-switch vertices and $D$, and hence construct $H$ from $G$. The same bound holds for the final “unplanting” $\triangle$-switch. The other three $\triangle$-switches in the simulation path match those in subcase (IXa), and we proved above that there are at most $4d_{1}M+2d_{1}^{4}+4d_{1}^{3}$ choices for $(G,H)$ if $(X,Y)$ is one of these three $\triangle$-switches. This gives at most
    $4d_{1}M+2d_{1}^{4}+4d_{1}^{3}+2(4d_{1}^{2}M)=8d_{1}^{2}M+4d_{1}M+2d_{1}^{4}+4d_{1}^{3}$
    choices for $(G,H)$ when $(X,Y)$ is part of a simulation path in subcase (IXc).
    Hence overall, there are at most
    $24d_{1}^{2}M+16d_{1}M+4d_{1}^{4}+8d_{1}^{3}+24d_{1}$
    choices for $(G,H)$ if $(X,Y)$ is part of a simulation path in Case (IX).
Adding up contributions from every case, we see that a $\triangle$-switch $(X,Y)$ can belong to the simulation path of at most
$528d_{1}^{2}M+32d_{1}M+4d_{1}^{4}+32d_{1}^{3}+14d_{1}^{2}+72d_{1}+$ 5
pairs $(G,G^{\prime})$. We can simplify this using the assumption that $d_{1}\geq 3$, and the fact that $M\geq 2d_{1}$ (as can be seen by considering a star $K_{1,d_{1}}$). This gives an upper bound of
$\left(28+\frac{32}{3}\right)d_{1}^{2}M+\left(4+\frac{32}{3}+\frac{14}{9}+\frac{72}{27}+\frac{5}{81}\right)d_{1}^{4}<20d_{1}^{2}\,(2M+d_{1}^{2}).$
This completes the proof. ∎
We can now prove Theorem 1.3.
Proof of Theorem 1.3.
Suppose that the switch chain is rapidly mixing on $\mathcal{G}_{n,\boldsymbol{d}}$ for all sequences $\boldsymbol{d}\in\mathcal{D}$. By Theorem 1.1, we can assume that the same is true for the modified Metropolis switch chain $\mathcal{M}^{\prime}$. That is, there exists a polynomial $p(n)$ such that if $\boldsymbol{d}\in\mathcal{D}$ and $\boldsymbol{d}$ has length $n$ then the mixing time of $\mathcal{M}^{\prime}$ on $\mathcal{G}_{n,\boldsymbol{d}}$ is at most $p(n)$. Using results of Sinclair [57], Guruswami [32, Theorem 4.9] proved that there exists a set $\Gamma^{\prime}$ of canonical paths for $\mathcal{M}^{\prime}$ and a polynomial $q(n)$ such that $\rho(\Gamma^{\prime})\leq q(n)$ for all $\boldsymbol{d}\in\mathcal{D}$ with length $n$.
Now we calculate the other parameters needed to apply Theorem 1.4. By Theorem 2.1, the maximum simulation path length is $\ell(\Sigma)=5$. Both chains have the same stationary distribution, namely the distribution which assigns to $G\in\mathcal{G}_{n,\boldsymbol{d}}$ the probability $\lambda^{\min\{t(G),\nu\}}/\hat{Z}_{\lambda}(\boldsymbol{d})$, where $\hat{Z}_{\lambda}(\boldsymbol{d})$ is the normalising factor. Hence the stationary ratio is given by $R(\mathcal{M},\mathcal{M}^{\prime})=1$. Similarly, the transition probability $P(G,H)$ is given by
$P(G,H)=\frac{1}{3\,a(\boldsymbol{d})}\,\min\{\lambda^{\min\{t(H),\nu\}-\min\{t(G),\nu\}},1\}$
whenever $G\neq H$ and $(G,H)$ is a switch (for $\mathcal{M}^{\prime}$) or a $\triangle$-switch (for $\mathcal{M}$). Therefore the simulation gap satisfies
$D(\mathcal{M},\mathcal{M}^{\prime})=\max_{\begin{subarray}{c}uv\in E(\mathcal{M})\\ zw\in E(\mathcal{M}^{\prime})\end{subarray}}\frac{\min\{\lambda^{\min\{t(z),\nu\}},\,\lambda^{\min\{t(w),\nu\}}\}}{\min\{\lambda^{\min\{t(u),\nu\}},\,\lambda^{\min\{t(v),\nu\}}\}}\leq\lambda^{\nu}\leq n^{\alpha},$
using the assumptions of the theorem and (12) for the final inequality. Hence it follows from Theorem 1.4 that there exists a set $\Gamma$ of canonical paths for $\mathcal{M}$ with congestion
$\bar{\rho}(\mathcal{M})\leq D(\mathcal{M},\mathcal{M}^{\prime})\,R(\mathcal{M},\mathcal{M}^{\prime})\ell(\Sigma)\,B(\Sigma)\,\bar{\rho}(\Gamma^{\prime})\leq 100n^{\alpha}\,d_{1}^{2}\,(2M+d_{1}^{2})\,q(n).$
Arguing as in the proof of Theorem 1.1 we can conclude that the mixing time $\tau_{\lambda}(\varepsilon)$ of the modified $\triangle$-switch chain is bounded above by
$\tau_{\lambda}(\varepsilon)\leq 100n^{\alpha}\,d_{1}^{2}\,(2M+d_{1}^{2})\,q(n)\,\big{(}M\log M+\alpha\log n+\log(\varepsilon^{-1})\big{)}.$
Thus, the modified Metropolis $\triangle$-switch chain with parameter $\lambda$ is rapidly mixing for all $\boldsymbol{d}\in\mathcal{D}$, under the assumption that $\lambda\mu\leq\log^{\alpha}n$ for some $\alpha\in(0,1)$. ∎
The assumption that $\lambda\mu\leq\log^{\alpha}n$ is fairly restrictive, but allows any constant value of $\lambda$ when $d_{1}=O(1)$. To go beyond this assumption would require a different approach to analysing the Markov chain.
3.1 Slow mixing
The upper bound on $\lambda\mu$ required for Theorem 1.3 is largely a consequence of the Metropolisation. Consider a $d$-regular graph $G$ which consists of $k$ disjoint copies of $K_{d+1}$. (Note that these graphs are central to the irreducibility proof in [16].) Then $G$ contains no paths of length 4, so no $\triangle^{+}$-switch is possible. Any $\triangle^{-}$-switch must take one edge from each of two disjoint copies of $K_{d+1}$, and hence will destroy $2(d-1)$ triangles. So any change to $G$ will occur only with probability $\lambda^{-2(d-1)}$. Suppose that $\lambda>1$ is any constant and $d=\gamma\log n$, where $\gamma\to\infty$ as $n\to\infty$. Then, with high probability, no change will occur within $(1/\lambda)^{d}=n^{\gamma\log\lambda}$ steps of the chain, which is superpolynomial.
There are exponentially few $d$-regular graphs of this form, as a fraction of the set of all $d$-regular graphs on $n=kd$ vertices, but each such graph gives a potential bottleneck for mixing. The small proportion suggests a solution by “ignoring” these graphs in some way. Unfortunately, the difficulty is not restricted to these graphs. Suppose we have only a single $K_{d+1}$ component. In order to break the $K_{d+1}$, we require the two vertices of a non-edge in the remainder of the graph to have many common neighbours, at least $d/4$, say. Then a $\triangle^{-}$-switch could create at most $d/2$ triangles, which is not enough to offset the $d-1$ triangles that have been destroyed. If there are $o(n)$ triangles in the current graph, then we can bound the probability of such a non-edge appearing in the graph by $O(n^{d/4+2}(d/n)^{d/2})=o(n^{-d/5})$, which is exponentially small if $d\to\infty$ with $n$. Thus all $\triangle^{-}$-switch’s involving the $K_{d+1}$ will have exponentially small probability of acceptance, and hence mixing will take exponential time.
Of course, the proportion of $d$-regular graphs with $n$ vertices and at least one $K_{d+1}$ component is still small. But, if $t(G)=\Omega(n)$, it is shown in [33] that $G$ contains many dense clique-like structures, called pseudocliques, which would require many steps to disassemble. Therefore, it seems unlikely that the $\triangle$-switch chain could generate regular graphs with $\Omega(n)$ triangles in polynomial time, except possibly for small constant $d$. Nevertheless, the modified $\triangle$-switch chain may well be rapidly mixing for a wider range of $\mu$ and $\lambda$ than we are able to prove here.
4 The asymptotic distribution of $t(G)$
The aim of this section is to analyse the distribution of the number of triangles in a random graph in $\mathcal{G}_{n,\boldsymbol{d}}$, both under the uniform distribution and the Gibbs distribution $\pi_{\lambda}$. We do this to justify the use of the cut-off $\nu$ in the modified Metropolis switch chain and the modified Metropolis $\triangle$-switch chain, see Remark 4, though the results of this section may also be of independent interest.
We denote the Poisson distribution with mean $\beta$ by $\textrm{Pois}(\beta)$ and, if the random variable $X$ has this distribution, we write in standard notation $X\thicksim_{d}\textrm{Pois}(\beta)$ or $\mathcal{L}(X)=\textrm{Pois}(\beta)$. We wish to determine the approximate distribution of $t(G)$ for large $n$, under the uniform distribution on $\mathcal{G}_{n,\boldsymbol{d}}$. We first give an appropriate definition of distributional approximation in this context, which has wider applicability. We note that the traditional approach to distributional approximation, particularly in Statistics, has been through limit distributions, usually with little or no error estimation.
4.1 Distributional approximation
Let $\mathcal{X}_{n}$ be a sequence of discrete state spaces, and let $X_{n},X^{\prime}_{n}$ be random variables on $\mathcal{X}_{n}$ such that $\mathcal{L}(X_{n})=p_{n},\mathcal{L}(X^{\prime}_{n})=p^{\prime}_{n}$. We will say that $X_{n}$ approximates $X^{\prime}_{n}$ asymptotically, which we denote by $X^{\prime}_{n}\thickapprox_{d}X_{n}$, if $p^{\prime}_{n}(A)=p_{n}(A)+o(1)$ for all $A\subseteq\Sigma$ as $n\to\infty$, That is, $X^{\prime}_{n}\thickapprox_{d}X_{n}$ if and only if
$\max_{A\subseteq\mathcal{X}}\big{|}\Pr(X^{\prime}_{n}\in A)-\Pr(X_{n}\in A)\big{|}=o(1).$
This involves a known metric on distributions, the total variation distance, $\mathrm{d}_{\mathrm{TV}}(p^{\prime}_{n},p_{n})$, as defined in (4). So $X^{\prime}_{n}\thickapprox_{d}X_{n}$ (and vice versa) if the total variation distance
$\mathrm{d}_{\mathrm{TV}}(p^{\prime}_{n},p_{n})=\tfrac{1}{2}\sum_{x\in\mathcal{X}_{n}}|p^{\prime}_{n}(x)-p_{n}(x)|=o(1)\mbox{\,\ as\ \,}n\to\infty.$
This relation is symmetric, reflexive and transitive, since $\mathrm{d}_{\mathrm{TV}}$ is a metric. However, transitivity holds only over a finite number of applications, since we must preserve the $o(1)$ term. So this is only a local equivalence relation on sequences $(\mathcal{X}_{n},p_{n})$.
If $X_{n}$ has a known distribution, $\textrm{Pois}(\beta)$ say, we also write $X^{\prime}_{n}\thickapprox_{d}\textrm{Pois}(\beta)$. This notation for distributional approximation is taken from [55], but was used only informally there.
This definition is not new. Janson [38] used the term “asymptotic equivalence” for this relationship between sequences of random variables, using the notation $(X_{n})\cong(X^{\prime}_{n})$. However, the focus in [38] is on the case where both $X^{\prime}_{n}$ and $X_{n}$ are unknown distributions, and the equivalence indicates a stronger relationship than contiguity. There the relationship between $X_{n}$ and $X^{\prime}_{n}$ is symmetrical. Here we assume $X_{n}$ has a known distribution which approximates the unknown distribution of $X^{\prime}_{n}$, so the relationship is not symmetrical. Also, as we have noted above, this is not a true equivalence relation. Therefore, we prefer the above terminology and notation.
Note that requiring only $p^{\prime}_{n}(x)=p_{n}(x)+o(1)$ for all $x\in\mathcal{X}_{n}$ is not a useful notion of approximation. For example, suppose $\mathcal{X}_{n}=[2n]$, $p_{n}(x)=1/n$ ($x\in[n]$), $p^{\prime}_{n}(x)=1/n$, ($x\in[2n]\setminus[n]$). Then $X_{n}$ approximates $X^{\prime}_{n}$ in this weaker sense, since $p^{\prime}_{n}(x)=p_{n}(x)\pm 1/n$. However, $\mathrm{d}_{\mathrm{TV}}(p_{n},p^{\prime}_{n})=1$ for all $n$, so this does not seem a reasonable approximation.
Though we will only use it in the discrete case, the approach to distributional approximation outlined above can be extended to arbitrary probability measures, as is done in [38]. It was initially developed for normal approximations [60], but later for other distributions (e.g. [13, 54, 55]). The methods used have been elementary (e.g. [56]), coupling arguments (e.g. [54]), Stein’s method [60, 8] and semigroup methods (e.g. [45, 59]). The aim has been to quantify errors in limit distributions in Statistics, or to quantify rates of convergence to these limit distributions, rather than to quantify asymptotic approximations.
We will use only elementary methods, but with an alternative criterion, which we now show is equivalent to the definition above.
Lemma 4.1.
$X_{n}\thickapprox_{d}X^{\prime}_{n}$ if and only if there is a set $S_{n}\subset\mathcal{X}_{n}$ such that, as $n\to\infty$, $p^{\prime}_{n}(x)=(1+o(1))p_{n}(x)$ for all $x\notin S_{n}$ and $\max\{p_{n}^{\prime}(S_{n}),p_{n}(S_{n})\}=o(1)$.
Proof.
If $S_{n}$ exists then there exist sequences $0<\delta_{n},\epsilon_{n}=o(1)$ so that $|p^{\prime}_{n}(x)-p_{n}(x)|<\delta_{n}p_{n}(x)$ for $x\notin S_{n}$ and $\max\{p_{n}^{\prime}(S_{n}),p_{n}(S_{n})\}<\epsilon_{n}$. Hence
$\mathrm{d}_{\mathrm{TV}}(p_{n},p^{\prime}_{n})<\epsilon_{n}+\delta_{n}\sum_{x\notin S_{n}}p_{n}(x)<\epsilon_{n}+\delta_{n}=o(1),$
so $X_{n}\thickapprox_{d}X^{\prime}_{n}$.
Conversely, if $X_{n}\thickapprox_{d}X^{\prime}_{n}$ then there is a sequence $0<\eta_{n}=o(1)$ such that $\mathrm{d}_{\mathrm{TV}}(p_{n},p^{\prime}_{n})<\eta_{n}$. Let $\epsilon_{n}=\sqrt{\eta_{n}}=o(1)$ and define
$S_{n}=\left\{x\in\mathcal{X}_{n}:\frac{|p^{\prime}_{n}(x)-p_{n}(x)|}{\max\{p_{n}(x),p^{\prime}_{n}(x)\}}\geq\epsilon_{n}\right\}.$
Then
$\epsilon_{n}^{2}>\mathrm{d}_{\mathrm{TV}}(p_{n},p^{\prime}_{n})\geq\epsilon_{n}\sum_{x\in S_{n}}\max\{p_{n}(x),p^{\prime}_{n}(x)\}\geq\epsilon_{n}\max\{p_{n}(S_{n}),p^{\prime}_{n}(S_{n})\},$
so $\max\{p_{n}(S_{n}),p^{\prime}_{n}(S_{n})\}<\epsilon_{n}=o(1)$.
If $x\notin S_{n}$, then we have $|p^{\prime}_{n}(x)-p_{n}(x)|<\epsilon_{n}\max\{p_{n}(x),p^{\prime}_{n}(x)\}$. By taking $n$ sufficiently large, we can assume that $\epsilon_{n}<\nicefrac{{1}}{{2}}$, and hence $p^{\prime}_{n}(x)\leq 2p_{n}(x)$ for all $x\not\in S_{n}$. For all such $n$ and all $x\not\in S_{n}$, we conclude that $p^{\prime}_{n}(x)/p_{n}(x)\in(1-2\epsilon_{n},1+2\epsilon_{n})$, and hence $p^{\prime}_{n}(x)\in(1\pm\delta_{n})p_{n}$ where $\delta_{n}=2\epsilon_{n}=o(1)$ as $n\to\infty$. This completes the proof. ∎
This formulation also gives explicit bounds on errors. In one direction, we can take $\eta_{n}=\epsilon_{n}+\delta_{n}$, while in the other we can take $\epsilon_{n}=\sqrt{\eta_{n}}$ and $\delta_{n}=2\sqrt{\eta_{n}}$.
We will apply Lemma 4.1 to prove the following result.
Theorem 4.1.
Suppose that $\boldsymbol{d}$ is a degree sequence with $d_{1}=o(n^{1/9})$ and $d_{n}\geq 1$, such that $M_{2}\geq M$. Let $\mathcal{X}_{n}=\{0,1,\ldots,M_{2}/6\}$.
  1. (i)
    Let $X_{n}^{\prime}$ be the number of triangles $t(G)$ when $G$ is chosen uniformly at random from $\mathcal{G}_{n,\boldsymbol{d}}$. Then $X_{n}^{\prime}\thickapprox_{d}\emph{$\textrm{Pois}(\mu)$}$ where $\mu=\mu(\boldsymbol{d})$.
  2. (ii)
    Further suppose that $\lambda>1$ is a constant with $d_{1}\log\lambda=o(\log n)$, and let $X_{n}^{\prime}$ be the number of triangles $t(G)$ when $G$ is chosen from $\mathcal{G}_{n,\boldsymbol{d}}$ using the Gibbs distribution $\pi_{\lambda}$. Then $X_{n}^{\prime}\thickapprox_{d}\emph{$\textrm{Pois}(\lambda\mu)$}$ where $\mu=\mu(\boldsymbol{d})$.
In particular, part (ii) implies that the mean number of triangles under $\pi_{\lambda}$ is $\lambda\mu$, compared with $\mu$ under the uniform distribution. Hence, in order to increase the number of triangles as far as possible, we should take $\lambda$ as large as possible, under the constraint $d_{1}\log\lambda=o(\log n)$.
To prove Theorem 4.1 we will take $S_{n}=\{G:t(G)>t_{0}\}$, for a suitably large $t_{0}$, to be determined. Thus $S_{n}$ is the tail of the distribution of $t(G)$. We divide the proof by analysing first “small” values of $t<t_{0}$ in Section 4.2 and then “large” values of $t>t_{0}$ in Section 4.2. Thus, in Section 4.2 we consider $G\notin S_{n}$, and in Section 4.3 we consider $G\in S_{n}$.
Before proceeding, we make some more remarks about distributional approximations. Note that $X^{\prime}_{n}\thickapprox_{d}X_{n}$ does not imply that $X^{\prime}_{n},X_{n}$ have exactly the same properties. Suppose $Y_{n}=h_{n}(X_{n}),Y^{\prime}_{n}=h_{n}(X^{\prime}_{n})$ for a function $h_{n}:\mathcal{X}_{n}\to\mathbb{C}$ (the complex numbers). Let $\|h_{n}\|=\max_{x\in\mathcal{X}_{n}}|h_{n}(x)|$. Then we can only claim $\mathbb{E}(Y^{\prime}_{n})\sim\mathbb{E}(Y_{n})$ if $h_{n}$ is uniformly bounded in $n$ using this norm. This follows from the (easily proved) inequality
$\big{|}\mathbb{E}[Y^{\prime}_{n}]-\mathbb{E}[Y_{n}]\big{|}\leq\|h_{n}\|\mathrm{d}_{\mathrm{TV}}(p_{n},p^{\prime}_{n}),$
with equality possible if $h_{n}(x)=\textrm{sign}\big{(}p^{\prime}_{n}(x)-p_{n}(x)\big{)}$, where $\textrm{sign}(x)\in\{-1,0,1\}$ (in the obvious way), so $\|h_{n}\|=1$ (unless $p^{\prime}_{n}=p_{n}$ and $\mathrm{d}_{\mathrm{TV}}(p^{\prime}_{n},p_{n})=0$). Indeed, this equality is another characterisation of $\mathrm{d}_{\mathrm{TV}}$. Note that the uniform boundedness requirement includes $X_{n}$ itself, so care is needed.
As an example, suppose $\mathcal{X}_{n}=[n]$, $p_{n}=C_{n}/i^{2}$ ($i\in[n]$), where $C_{n}\sim\pi^{2}/6$. (This example is related to the Cauchy distribution.) However, $\mathbb{E}[X_{n}]=\Omega(\log n)\to\infty$ as $n\to\infty$. The explanation is that $n$ is not a uniformly bounded function of $n$. This is a known pathology in the limit theory of distributions. In this example, $X^{\prime}_{n}$ converges pointwise to a fixed distribution, but $\lim_{n\to\infty}\mathbb{E}[X^{\prime}_{n}]$ does not even exist.
In the limit distribution approach to approximation, we must restrict attention to uniformly bounded functions. Here, if $h_{n}$ is not uniformly bounded, we have the additional option of estimating the contribution to the expectation from $S_{n}$, as we do below for Poisson approximations.
However, if $X^{\prime}_{n}\thickapprox_{d}X_{n}$ for real-valued random variables, then $X_{n}$ and $X^{\prime}_{n}$ do have the same characteristic function $\mathbb{E}(e^{itx})$ in the limit as $n\to\infty$, since $|e^{itx}|=1$ for all real $t$ and $x$. This could be used, for example, to prove asymptotic normality of the Poisson approximations below in the limit distribution sense.
Note that we cannot use this idea of approximation to compare discrete distributions with continuous distributions. Discrete points have measure zero in a continuous distribution, so we always have $\mathrm{d}_{\mathrm{TV}}(X_{n},X^{\prime}_{n})=1$. Thus, any definition of approximation in this context must use a weaker metric than $\mathrm{d}_{\mathrm{TV}}$. For real-valued random variables, the usual choice is the Kolmogorov metric $\max_{x\in\mathbb{R}}|F_{n}(x)-F^{\prime}_{n}(x)|$, where $F_{n},F^{\prime}_{n}$ are the respective distribution functions of $X_{n},X^{\prime}_{n}$. Thus, the asymptotic normal results for triangles proved in [24] are valid for larger values of $\mu(\boldsymbol{d})$ than the asymptotic Poisson results proved here, but they are much weaker approximations. In fact, the indirect method of proof used in [24] does not seem to allow for any quantification of errors.
4.2 Small number of triangles
In this section we will prove the following result.
Theorem 4.2.
Let $G$ be chosen uniformly from $\mathcal{G}_{n,\boldsymbol{d}}$ and let $\mu=\mu(\boldsymbol{d})$. Suppose that $M_{2}\geq M$ and that $t_{0}=\min\{\sqrt{M/d_{1}^{3}},\omega(n)d_{1}^{3}\}$ for some function $\omega(n)\to\infty$ as $n\to\infty$. Then $\Pr(t(G)=t)\sim e^{-\mu}\mu^{t}/t!$ for all $t=o(t_{0})$.
Our proof will be based on the switchings technique of McKay [48] and others, combined with the conditional probability estimates of Gao and Ohapkin [26]. The switchings technique has been applied in the analysis of the switch chain before, see for example [25, 30]. For reference, we paraphrase [26, Corollary 2] as follows.
Lemma 4.2.
([26, Corollary 2])  Let $H^{+}\in\mathcal{G}_{n,\boldsymbol{d}^{\prime}}$, where $\boldsymbol{d}^{\prime}<\boldsymbol{d}$, and let $M^{\prime}=\sum_{i=1}^{n}d^{\prime}_{i}=2|E(H^{+})|$. Then, if $G$ is chosen uniformly from $\mathcal{G}_{n,\boldsymbol{d}}$,
$\Pr(uv\in G\mid H^{+}\subset G)=\bigg{(}1+O\Big{(}\frac{d_{1}^{2}}{M-M^{\prime}}\Big{)}\bigg{)}\frac{(d_{u}-d^{\prime}_{u})(d_{v}-d^{\prime}_{u})}{M-M^{\prime}+(d_{u}-d^{\prime}_{u})(d_{v}-d^{\prime}_{u})}.$
In our application, $H^{+}$ will be a subgraph $H_{t}$ of $G$ containing exactly $t$ triangles with $t=o(M)$, which will imply that the error term in Lemma 4.2 is $O(d_{1}^{3}/M)$.
Remark 1.
In [26] there is also a subgraph $H^{-}$, edge-disjoint from $H^{+}$, which is conditioned on being absent from $G$. Since we make no use of this, we omit it. Also, we make the $1+o(1)$ term explicit in Lemma 4.2, since we iterate these estimates and hence exponentiate this term. We start with $H^{+}=\emptyset$, $\boldsymbol{d}^{\prime}=\mathbf{0}$, so $H^{+}$ remains small relative to $G$.
Remark 2.
Note that Lemma 4.2 gives essentially the unconditioned estimate on the graph which results from deleting the edges of $H^{+}$ from $G$, as might be expected using the configuration model for graphs with given degree sequences.
We apply these estimates with the switching method, as follows. For $t=0,\ldots,M_{2}/6$, let $\mathcal{N}_{t}=\{G\in\mathcal{G}_{n,\boldsymbol{d}}:t(G)=t\}$ be the set of graphs with exactly $t$ triangles, and let $N_{t}=|\mathcal{N}_{t}|$. We will use $G_{t}$ to denote an element of $\mathcal{N}_{t}$ chosen uniformly at random. Consider the set $S_{t}^{+}$ of $\triangle^{+}$-switches from graphs in $\mathcal{N}_{t}$ which create exactly one triangle and destroy none. Also define the set of the inverses of these $\triangle^{+}$-switches, namely the set $S_{t+1}^{-}$ of $\triangle^{-}$-switches from graphs in $\mathcal{N}_{t+1}$ which delete exactly one triangle and create none. Let $s_{t}^{+}=|S_{t}^{+}|$ and $s_{t}^{-}=|S_{t}^{-}|$. Then we have
$N_{t}\,\mathbb{E}[s_{t}^{+}\mid G_{t}]\,=\,N_{t+1}\,\mathbb{E}[s_{t+1}^{-}\mid G_{t+1}],$ (14)
from which we can estimate $N_{t+1}/N_{t}$. We note that McKay and co-authors have considerably generalised this switchings idea, for example in [35], but here we only require the simple version given above.
Recall that $\mu=\mu(\boldsymbol{d})=\frac{M_{2}^{3}}{6M^{3}}$. We first prove a claim made in Section 1.1 above.
Lemma 4.3.
Let $G$ be chosen uniformly from $\mathcal{G}_{n,\boldsymbol{d}}$. If $d_{1}^{2}=o(M)$ and $M_{2}\geq M$ then $\mathbb{E}[t(G)]\sim\mu.$
Proof.
Using Lemma 4.2, and calculations similar to cases (a) and (b) of Lemma 4.4 below, we have, for triangles $i,j,k$,
$\mathbb{E}[t(G)]=\Big{(}1+O\Big{(}\frac{d_{1}^{2}}{M}\Big{)}\Big{)}^{3}\,\sum_{\begin{subarray}{c}i,j,k=1\\ \text{distinct}\end{subarray}}^{n}\,\frac{d_{i}(d_{i}-1)d_{j}(d_{j}-1)d_{k}(d_{k}-1)}{6M^{3}}\sim\frac{M_{2}^{3}}{6M^{3}}=\mu,$ (15)
where the divisor 6 counts permutations of $i,j,k$ representing the same triangle. The exponent 3 on the error term in (15) results from three uses of Lemma 4.2 and replacing the three occurrences of $M+O(d_{1}^{2})$ with $M$. A relative error of $O(d_{1}^{2}/M_{2})$ arises from the terms where $i,j,k$ are not distinct: these terms arise in $M_{2}^{3}$ but are not present in the original sum. Therefore we need to assume $M_{2}=\Omega(M)$, but we make the stronger assumption $M_{2}\geq M$. ∎
We have $\mu(\boldsymbol{d})=\frac{M_{2}^{3}}{6M^{3}}\geq\nicefrac{{1}}{{6}}$ under our assumption that $M_{2}\geq M$, so $\mu=\Omega(1)$ as $n\to\infty$.
Note that the numerator in the summation in (15) is independent of the order in which the edges of the triangle are exposed, and depends only on its degree structure. This observation is true for all small subgraphs, and we use it below without further comment.
Lemma 4.4.
[…]
Proof.
We first give an estimate of $\mathbb{E}[s_{t}^{+}\mid G_{t}]$, and then show that it is asymptotically tight. A $\triangle^{+}$-switch corresponds to a path of length 5 in $G$, which we will denote by $P_{5}$, though we make no assumptions about other possible edges between its vertices. The $\triangle^{-}$-switch corresponding to this $P_{5}$ starts with a triangle on its 3 middle edges, and a disjoint edge between its end vertices. We will label these as illustrated below.
$i$$j$$k$$\ell$$m$$i$$j$$k$$\ell$$m$
Then our initial estimate is the number of $P_{5}$’s, where we will temporarily ignore other restrictions:
$\mathbb{E}[s_{t}^{+}\mid G_{t}]$ $\leq\big{(}1+O(d_{1}^{2}/M)\big{)}^{4}\,\sum_{\begin{subarray}{c}i,j,k,\ell,m=1\\ \text{distinct}\end{subarray}}^{n}\frac{d_{i}(d_{i}-1)\,d_{j}(d_{j}-1)\,d_{k}(d_{k}-1)\,d_{\ell}\,d_{m}}{M^{4}}$
$\leq\big{(}1+O(d_{1}^{2}/M)\big{)}\,\frac{M_{2}^{3}M^{2}}{M^{4}}=6M\mu,$ (16)
where the expression in the summation is obtained using Lemma 4.2 with $H_{t}=\emptyset$.
The estimate (16) is an upper bound because we have ignored conditions that should be imposed and which reduce the expectation. We will now refine this by considering each possible overestimate in turn. This will give us a lower bound on the expectation which matches the upper bound asymptotically for small enough $t$ and $d_{1}$.
  1. (a)
    First, we should have $i\neq j\neq k\neq\ell\neq m$. There are $\binom{5}{2}=10$ possible first-order equalities in (16), and we claim that each of these gives rise to a $1+O(d_{1}^{2}/M)$ error factor. Consider, for example if $i=j$. If the other vertices are distinct, then (16) becomes
    $\sum_{i,k,\ell,m=1}^{n}\frac{d_{i}^{2}(d_{i}-1)^{2}d_{k}(d_{k}-1)d_{\ell}d_{m}}{M^{4}}\leq\frac{d_{1}^{2}M_{2}^{2}M^{2}}{M^{4}}=\frac{d_{1}^{2}M\mu}{M_{2}}\leq d_{1}^{2}\mu.$
    Comparing this with (16) gives the claim. So overall the effect is no worse than a $(1+O(d_{1}^{2}/M))^{10}=1+O(d_{1}^{2}/M)$ factor. Note that higher-order equalities are automatically included, in fact over-counted, in these calculations.
  2. (b)
    We have used $M$ in the denominator in (16) rather than the $M-M^{\prime}$ from Lemma 4.2. In fact, $H_{t}$ is a set of $t$ triangles, so $M^{\prime}\leq 6t$, so the denominator in Lemma 4.2 is at most $M+d_{1}^{2}$, and at least $M-6t$. So using $M$ produces a $(1+O(t+d_{1}^{2})/M)^{4}=1+O(t+d_{1}^{2})/M$ error factor in (16).
  3. (c)
    We have not considered the effect of $H_{t}$ on the calculation of the quantity $M^{\prime}_{2}=\sum_{i=1}^{n}(d_{i}-d^{\prime}_{i})(d_{i}-d^{\prime}_{i}-1)$ which corresponds to $M_{2}$ on $G$. We have
    $M_{2}\geq M^{\prime}_{2}=\sum_{i=1}^{n}(d_{i}-d^{\prime}_{i})(d_{i}-d^{\prime}_{i}-1)\geq\sum_{i=1}^{n}d_{i}(d_{i}-1)-2d_{1}\sum_{i=1}^{n}d^{\prime}_{i}\geq M_{2}-12td_{1}.$
    So using $M_{2}$ produces a $(1+O(td_{1})/M_{2})^{3}=1+O(td_{1})/M$ error factor in (16).
  4. (d)
    We wish to ensure that almost all $P_{5}$’s do not contain an edge of $H_{t}$. We claim that these give rise to a $1+O(td_{1}^{2}/M)$ error factor. Consider, for example if $ij\in E(H_{t})$. Then (16) becomes
    $2\sum_{ij\in E(H_{t})}\sum_{k,\ell,m=1}^{n}\frac{(d_{i}-1)(d_{j}-1)d_{k}(d_{k}-1)d_{\ell}d_{m}}{M^{3}}$ $\leq\frac{2td_{1}^{2}M_{2}M^{2}}{M^{3}}=\frac{2td_{1}^{2}M_{2}^{3}M^{2}}{M^{3}M_{2}^{2}}$
    $=\frac{12td_{1}^{2}\mu M^{2}}{M_{2}^{2}}=O(td_{1}^{2}\mu).$
    Again, comparing this with (16) gives the claim. Since there are only four edges, the overall the effect is no worse than a factor $(1+O(td_{1}^{2}/M))^{4}=1+O(td_{1}^{2}/M)$.
    Since it avoids edges of $H_{t}$, the $\triangle^{+}$-switch move has probability at most $O(td_{1}^{2}\mu)/M$ of breaking any triangle.
  5. (e)
    We need to ensure that the edges $jk$ and $\ell m$ are absent in almost all $P_{5}$’s, so that they give a valid $\triangle^{+}$-switches. We claim that these give rise to a $1+O(d_{1}^{2}/M)$ error factor. Consider $jk$. The expected number of $P_{5}$’s in which it is present is at most
    $\sum_{i,j,k,\ell,m=1}^{n}\frac{d_{i}(d_{i}-1)d_{j}(d_{j}-1)(d_{j}-2)d_{k}(d_{k}-1)(d_{k}-2)d_{\ell}d_{m}}{M^{5}}\leq\frac{d_{1}^{2}M_{2}^{3}M^{2}}{M^{5}}=O(\mu d_{1}^{2}),$
    and comparison with (16) gives the claim. The calculation for $\ell m$ is similar.
  6. (f)
    We wish to ensure that the edges $jk$ and $\ell m$ are not edges of $H_{t}$. Otherwise, the $P_{5}$ will not give a valid $\triangle^{+}$-switch. If $jk\in H_{t}$, then (16) becomes
    $2\sum_{jk\in E(H_{t})}\sum_{i,\ell,m=1}^{n}$ $\frac{d_{i}(d_{i}-1)(d_{j}-2)(d_{k}-2)d_{\ell}d_{m}}{M^{4}}\leq\frac{2td_{1}^{2}M_{2}M^{2}}{M^{4}}=\frac{2td_{1}^{2}M_{2}}{M^{2}}$
    $=\frac{12td_{1}^{2}\mu}{M}=O(td_{1}^{2}\mu/M).$
    and, if $\ell m\in E(H_{t})$, (16) becomes
    $2\sum_{\ell m\in E(H_{t})}\sum_{i,j,k=1}^{n}$ $\frac{d_{i}(d_{i}-1)d_{j}(d_{j}-2)d_{k}(d_{k}-1)(d_{\ell}-1)(d_{m}-1)}{M^{4}}\leq\frac{2td_{1}^{2}M_{2}^{3}}{M^{4}}$
    $=\frac{12td_{1}^{2}\mu}{M}=O(td_{1}^{2}\mu/M).$
    Note that these events will not even occur asymptotically if $td_{1}^{2}\mu=o(M)$, since their expectation is then $o(1)$.
  7. (g)
    We must ensure that almost all $P_{5}$’s create only one triangle. The expected number in which the new edge $jk$ will create a second triangle $(hjk)$ is at most
    $\sum_{h,i,j,k,\ell,m=1}^{n}$ $\frac{d_{h}(d_{h}-1)d_{i}(d_{i}-1)d_{j}(d_{j}-1)(d_{j}-2)d_{k}(d_{k}-1)(d_{k}-2)d_{\ell}d_{m}}{M^{6}}$
    $\leq\frac{d_{1}^{2}M_{2}^{4}M^{2}}{M^{6}}=\frac{6d_{1}^{2}\mu M_{2}}{M}=O(\mu d_{1}^{3}).$
    The expected number of $P_{5}$’s in which the new edge $\ell m$ will create any triangle $(h\ell m)$ is at most
    $\sum_{h,i,j,k,\ell,m=1}^{n}$ $\frac{d_{i}(d_{i}-1)d_{j}(d_{j}-1)d_{k}(d_{k}-1)d_{\ell}(d_{\ell}-1)d_{h}(d_{h}-1)d_{m}(d_{m}-1)}{M^{6}}$
    $=\frac{M_{2}^{6}}{M^{6}}=\mu^{2}=O(\mu d_{1}^{3}).$
    Combining these two cases gives an error factor $1+O(d_{1}^{3}/M)$.
The effect of combining all these cases gives an overall error factor which we may bound by $1+O(d_{1}^{2}(d_{1}+t)/M)=1+O(td_{1}^{3}/M)$, the critical cases arising from (d) and (g). Note also that (d) implies that almost all valid $\triangle^{+}$-switches give rise to a triangle in $G_{t+1}$ which is edge-disjoint from $H_{t}$. ∎
We now consider the $\triangle^{-}$-switches reversing these $\triangle^{+}$-switches.
Lemma 4.5.
Suppose that $M_{2}\geq M$. If $td_{1}^{2}=o(M)$, then
$\mathbb{E}[s_{t+1}^{-}\mid G_{t}]\,=\,(1+O(td_{1}^{3}/M))6(t+1)M.$
Proof.
We must choose a triangle $(ijk)$ and a labelling of it as $i,j,k$, which can be done in $6$ ways. Then, from the proof of Lemma 4.4 we may assume that the triangle $(ijk)$ is edge-disjoint from any other since, from case (d) in the proof of Lemma 4.4, this will be true for all but a proportion $O(td_{1}^{2}/M)$. So deleting $jk$ will delete exactly one triangle. Note that we could show that $(ijk)$ will be vertex-disjoint from $H_{t}$, but edge-disjointness is sufficient here.
Deleting $\ell m$ will delete no triangles if $\ell m\notin E(H_{t})$. So there are at most $M$, and at least $M-2|E(H_{t})|$ choices for $\ell m$. So there are $(1-O(t/M))M$ choices. If $\ell j\in E(H_{t})$ or $mk\in E(H_{t})$, the $\triangle^{-}$-switch will be invalid, but again case (d) in the proof of Lemma 4.5 implies that only a proportion $O(td_{1}^{2}/M)$ of $\triangle^{-}$-switches will $G_{t+1}$ can be invalid for this reason. Otherwise, the expected number of triangles $(\ell jh)$ created by inserting $\ell j$ will be at most
$\sum_{h=1}^{n}\frac{d_{h}(d_{h}-1)(d_{j}-2)(d_{\ell}-1)}{M^{2}}\leq\frac{d_{1}^{2}M_{2}}{M^{2}}\leq\frac{d_{1}^{3}}{M}.$
So the error factor in $\mathbb{E}[s_{t+1}^{-}\mid G_{t}]$ is $1+O(d_{1}^{3}/M)$. The same calculation for $mk$ gives the same error, so the total effect will remain $1+O(d_{1}^{3}/M)$.
Combining the error estimates we have an overall $1+O(td_{1}^{3}/M)$ factor, completing the proof. ∎
We can now prove Theorem 4.2.
Proof of Theorem 4.2.
It follows from Lemmas 4.4 and 4.5 and (14) that
$\frac{N_{t+1}}{N_{t}}=(1+O(td_{1}^{3}/M))\frac{\mu}{t+1}.$
Iterating this through $t,t-1,\ldots,1$,
$\frac{\Pr(t(G)=t)}{\Pr(t(G)=0)}=\frac{N_{t}}{N_{0}}=(1+O(td_{1}^{3}/M))^{t}\frac{\mu^{t}}{t!}=(1+O(t^{2}d_{1}^{3}/M))\frac{\mu^{t}}{t!}\sim\frac{\mu^{t}}{t!},$ (17)
if $t^{2}d_{1}^{3}=o(M)$, which is $t=o(\sqrt{M/d_{1}^{3}}$), so is valid for $t\leq t_{0}$. Summing (17)
$\frac{\Pr(t\leq t_{0})}{\Pr(t=0)}\sim\sum_{t=0}^{t_{0}}\frac{\mu^{t}}{t!}=e^{\mu}-\sum_{t=t_{0}+1}^{\infty}\frac{\mu^{t}}{t!}.$ (18)
Now $d_{1}^{3}=o(t_{0})$ implies $\mu=o(t_{0})$, so the term on the right is
$\sum_{t=t_{0}+1}^{\infty}\frac{\mu^{t}}{t!}<\sum_{t=t_{0}+1}^{\infty}\left(\frac{o(t_{0})}{t_{0}}\right)^{t}\sim o(1)^{t_{0}}=o(1).$
In Section 4.3 below, we show that $\Pr(t\leq t_{0})=1-o(1)$. So, since $\mu=\Omega(1)$, (18) gives
$\frac{\Pr(t\leq t_{0})}{\Pr(t=0)}=\frac{1-o(1)}{\Pr(t=0)}\sim e^{\mu}-o(1)=(1-o(1))e^{\mu}.$ (19)
From (19), $\Pr(t=0)\sim e^{-\mu}$, and so (17) gives $\Pr(t(G)=t)\sim e^{-\mu}\mu^{t}/t!$, and so $t(G)$ is approximately Poiss($\mu$). ∎
We state a useful special case.
Corollary 4.1.
Let $G$ be chosen uniformly from $\mathcal{G}_{n,\boldsymbol{d}}$ and let $\mu=\mu(\boldsymbol{d})$. Suppose that $M_{2}\geq M$. If $d_{1}=o(n^{1/9})$ and $d_{n}\geq 1$ then $\Pr(t(G)=t)\sim e^{-\mu}\mu^{t}/t!$ for all $t=o(n^{1/3})$.
Proof.
Taking the $\omega(n)$ function in Theorem 4.2 to equal $n^{1/3}/d_{1}^{3}$, the given assumptions imply that $\omega(n)\to\infty$ and that the function $t_{0}$ defined in the statement of Theorem 4.2 satisfies $t_{0}\geq n^{1/3}$. The result now follows follows from Theorem 4.2. ∎
Remark 3.
Gao [24] has shown a normal limit distribution for $t(G)$ for $d$-regular graphs as $n\to\infty$, in the form $(t-\mu)/\sqrt{\mu}\stackrel{{\scriptstyle d}}{{\to}}\mathrm{N}(0,1)$, provided $d\to\infty$ and $d=O(\sqrt{n})$. This is a limit for $\textrm{Pois}(\mu)$ as $\mu\to\infty$, but whether Poisson approximation holds for such large $d$ is an open question.
4.3 Large number of triangles
Here we will establish much weaker bounds, since our goal is only to show that the tail of the distribution of $t(G)$ has very small probability in the equilibrium distribution, provided $\lambda$ and $d_{1}$ are small enough. Specifically, we prove the following. Our argument follows the approach of the general-purpose result of Hasheminezhad and McKay [35, Corollary 1], but to make our proof self-contained we give the details here.
Theorem 4.3.
Let $t_{0}=\Omega(n^{1/3})$ and suppose that $d_{1}=o(n^{1/9})$. Let $\lambda\geq 1$ be constant. If $\lambda>1$ then we also assume that $d_{1}\log\lambda=o(\log n)$. If $G$ is chosen from $\mathcal{G}_{n,\boldsymbol{d}}$ using the distribution $\pi_{\lambda}$ then
${\Pr_{\pi_{\lambda}}}(t\geq t_{0})=e^{-\Omega(n^{1/4})}.$
Here the implicit constant inside the $\Omega(\cdot)$ term is positive.
Proof.
Given $G\in\mathcal{N}_{t}$, let $s^{-}(G)$ be the number of valid $\triangle^{-}$-switches from $G\in\mathcal{G}_{i}$ which create no triangle. A lower bound on $s^{-}(G)$ is given by taking any triangle and any edge at distance at least 3 from the triangle. There are at least $M-6(d_{1}-1)^{3}$ choices for this directed edge. If $d_{1}=o(M^{1/3})$, this is $(1-o(1))M$. So $s^{-}(G)\geq(6-o(1))tM$ since we must choose a triangle edge and a direction, and $s^{-}(\mathcal{N}_{t})=\sum_{G\in\mathcal{N}_{t}}s^{-}(G)\geq(6-o(1))tMN_{t}$. No triangles are created by any of these $\triangle^{-}$-switches, but more than one triangle may be destroyed by removal of the switch edges. So we destroy at least 1 and at most $2(d_{1}-1)$ triangles. These bounds are tight, since every edge might lie in a $(d_{1}+1)$-clique, as is the case for graphs comprising $n/(d_{1}+1)$ cliques of size $(d_{1}+1)$.
So all the $\triangle^{-}$-switches from $G\in\mathcal{N}_{t}$ lead to a graph in $\mathcal{N}^{\prime}_{t}=\bigcup_{i=1}^{2(d_{1}-1)}\mathcal{N}_{t-i}$. Suppose a proportion $\gamma_{i}$ lead to $\mathcal{N}_{t-i}$, so $\sum_{i=1}^{2(d_{1}-1)}\gamma_{i}=1$, and note that hence
$\sum_{i=1}^{2(d_{1}-1)}\gamma_{i}N_{t-i}\leq\max\{N_{t-i}:1\leq i\leq 2(d_{1}-1)\}=N_{t^{\prime}}$
for some $t-2(d_{1}-1)\leq t^{\prime}\leq t-1$. Observe that we will have $t^{\prime}=t-2(d_{1}-1)$ if $G$ is $d_{1}$-regular and $t=nd_{1}(d_{1}-1)/6$, so the lower bound can be tight. Moreover, van der Hoorn et al. [33] show that almost all $d_{1}$-regular graphs have almost all their triangles in large cliques for $t(G)=\Omega(n)$, so the analysis we give here may be difficult to improve significantly.
An upper bound on the number $s^{+}(G^{\prime})$ of valid $\triangle^{+}$-switches from any $G^{\prime}\in\mathcal{N}^{\prime}_{t}$ is the total number of possible switches, valid or not. There are at most $\sum_{i=1}^{n}2\binom{d_{i}}{2}(d_{1}-1)^{2}=M_{2}(d_{1}-1)^{2}$ possibilities, so there are at most $M_{2}(d_{1}-1)^{2}N_{t-i}$ valid $\triangle^{+}$-switches from $\mathcal{N}_{t-i}$ to $\mathcal{N}_{t}$. Consequently, we must have
$\frac{\gamma_{i}N_{t}}{N_{t^{\prime}}}\leq\frac{\gamma_{i}N_{t}}{N_{t-i}}\leq(1+o(1))\frac{M_{2}(d_{1}-1)^{2}}{6tM},$
and summing over $i$ gives
$\frac{N_{t}}{N_{t^{\prime}}}\leq(1+o(1))\frac{M_{2}(d_{1}-1)^{2}}{6tM}\leq\frac{\hat{\mu}}{t},$
where $\hat{\mu}=(d_{1}-1)^{3}/6\geq M_{2}^{3}/M^{3}=\mu$, since $M_{2}\leq M(d_{1}-1)$. Note that the final inequality is an equality if $G$ is $d_{1}$-regular.
We can iterate this to give
$\frac{N_{t}}{N_{0}}\leq\frac{\hat{\mu}^{\ell}}{tt^{\prime}\cdots t^{(\ell-1)}},$
where $t^{(j-1)}-t^{(j)}\leq 2d_{1}-2$ for $0<j<\ell$, $t^{(0)}=t$, $t^{(\ell)}=0$. The weakest bound results when $t^{(j-1)}-t^{(j)}=2d_{1}-2$ and $\ell=t/(2d_{1}-2)$, giving
$\frac{N_{t}}{N_{0}}\leq\frac{\hat{\mu}^{t/(2d_{1}-2)}}{(2d_{1}-2)^{t/(2d_{1}-2)}(t/(2d_{1}-2))!}=\frac{(\hat{\mu}/(2d_{1}-2))^{t/(2d_{1}-2)}}{(t/(2d_{1}-2))!}<\bigg{(}\frac{e\hat{\mu}}{t}\bigg{)}^{t/(2d_{1}-2)},$
using the bound $(t/(2d_{1}-2))!>(e^{-1}t/(2d_{1}-2))^{t/(2d_{1}-2)}$ from Stirling’s approximation. It follows that, when $G$ is chosen uniformly at random from $\mathcal{G}_{n,\boldsymbol{d}}$,
$\Pr(t(G)=t)=\frac{N_{t}}{|\mathcal{G}_{n,\boldsymbol{d}}|}<\bigg{(}\frac{e\hat{\mu}}{t}\bigg{)}^{t/(2d_{1}-2)}.$
If $\lambda>1$ then we have assumed that $d_{1}\log\lambda=o(\log n)$, and hence $\lambda^{2d_{1}-2}\hat{\mu}=o(t_{0})$. It follows for $\lambda\geq 1$ that $\lambda^{2d_{1}-2}\hat{\mu}\leq t_{0}/e^{2}$. Then under the Gibbs distribution $\pi_{\lambda}$,
${\textstyle{\Pr_{\pi_{\lambda}}}}(t\geq t_{0})\ =\sum_{t=t_{0}}^{M_{2}/6}\frac{\lambda^{t}\,N_{t}}{Z_{\lambda}(\boldsymbol{d})}$ $\leq\sum_{t=t_{0}}^{M_{2}/6}\frac{\lambda^{t}\,N_{t}}{|\mathcal{G}_{n,\boldsymbol{d}}|}$
$<$ $\ \sum_{t=t_{0}}^{M_{2}/6}\lambda^{t}\bigg{(}\frac{e\hat{\mu}}{t}\bigg{)}^{t/(2d_{1}-2)}<\ \sum_{t=t_{0}}^{\infty}\bigg{(}\frac{e\lambda^{2d_{1}-2}\hat{\mu}}{t}\bigg{)}^{t/(2d_{1}-2)}$
$<$ $\ \sum_{t=t_{0}}^{\infty}e^{-t/(2d_{1}-2)}\ =\ e^{-t_{0}/(2d_{1}-2)}/\big{(}1-e^{-1/(2d_{1}-2)}\big{)}$
$<$ $\ 2d_{1}e^{-t_{0}/(2d_{1}-2)}\ =\ e^{-\Omega(n^{1/4})},$
using the inequality $1-x<e^{-x}$ for $x>0$. This concludes the proof. ∎
We can now prove Theorem 4.1.
Proof of Theorem 4.1.
Part (i) follows directly from Lemma 4.1 using Corollary 4.1 and Theorem 4.3 with $\lambda=1$.
For (ii), we must look at the distribution of $t(G)$ when $G$ is chosen from $\mathcal{G}_{n,\boldsymbol{d}}$ using the Gibbs distribution $\pi_{\lambda}$. Let $\pi^{\prime}$ denote the uniform distribution on $\mathcal{G}_{n,\boldsymbol{d}}$. Using Corollary 4.1 and Theorem 4.3,
$\mathbb{E}_{\mathrm{\pi^{\prime}}}[\lambda^{t}]\sim\sum_{k\geq 0}\lambda^{k}e^{-\mu}\frac{\mu^{k}}{k!}\ \sim e^{(\lambda-1)\mu}\,\sum_{k\geq 0}\frac{(\mu\lambda)^{k}}{k!}e^{-\mu\lambda}=e^{(\lambda-1)\mu}.$
Hence, for the Gibbs distribution $\pi_{\lambda}$, if $s=o(n^{1/3})$ then
${\textstyle{{\Pr}_{\pi_{\lambda}}}}(t(G)=s)=\frac{\lambda^{s}\Pr_{\mathrm{\pi^{\prime}}}(t(G)=s)}{\mathbb{E}_{\mathrm{\pi^{\prime}}}[\lambda^{t}]}\sim\frac{\lambda^{s}e^{-\mu}\mu^{s}/s!}{e^{(\lambda-1)\mu}}=\frac{e^{-\lambda\mu}(\lambda\mu)^{s}}{s!}.$
Combining this with Theorem 4.3 shows that $t(G)\thickapprox_{d}\textrm{Pois}(\lambda\mu)$ under $\pi_{\lambda}$, completing the proof of part (ii). ∎
Remark 4.
If the assumptions of Theorem 4.3(ii) hold and $G$ is a random element of $\mathcal{G}_{n,\boldsymbol{d}}$ drawn from the distribution $\pi_{\lambda}$, the probability that $t(G)>\nu$ is $o(1/q(n))$ for any polynomial $q(n)$. Hence, after a “burn-in” period, if the modified Metropolis switch chain (Figure 3) is close to equilibrium and we observe it for polynomially many steps, it should be indistinguisable from the (unmodified) Metropolis switch chain (Figure 2). The same is true for the modified Metropolis $\triangle$-switch chain (Figure 4), compared with the unmodified Metropolis $\triangle$-switch switch chain. (The transition procedure of the unmodified Metropolis $\triangle$-switch chain is obtained from Figure 4 by replacing $\min\{t(H),\nu\}$ by $t(H)$ and replacing $\min\{t(G),\nu\}$ by $t(G)$.)
4.4 Assumptions on the degree sequence
In our asymptotic results we made the assumption that $M_{2}\geq M$. We will now prove that this is equivalent to having average degree $\bar{d}\geq 2$. For our Markov chain analysis the condition $\bar{d}\geq 2$ is automatically satisfied, since we assume that $d_{n}\geq 3$ in Theorem 1.3.
Lemma 4.6.
$M_{2}\geq M$ if and only if $\bar{d}\geq 2$.
Proof.
We have
$0\leq\sum_{i=1}^{n}(d_{i}-\bar{d})^{2}=\sum_{i=1}^{n}d_{i}^{2}-n\bar{d}^{2}=M_{2}+M-M\bar{d}=M_{2}-M(\bar{d}-1),$
so $M_{2}\geq M$ if $\bar{d}\geq 2$. For the converse, suppose that $M_{2}<M$. Let $n_{j}$ be the number of vertices with degree $j$, for all $j\geq 1$. Then
$M_{2}-M=-n_{1}+\sum_{j\geq 3}j(j-2)n_{j}<0,$
and
$M=n_{1}+2n_{2}+\sum_{j\geq 3}jn_{j}\leq n_{1}+2n_{2}+\sum_{j\geq 3}j(j-2)n_{j}<2n_{1}+2n_{2}\leq 2n.$
So $\bar{d}=M/n<2$, completing the proof. ∎
The excluded graphs are rather uninteresting. They have a preponderance of components which are paths, as we see from the following.
Lemma 4.7.
$M_{2}\geq M$ for any connected graph $G$ which is not a path.
Proof.
If $G$ is not a path or a cycle, then $d_{1}\geq 3$. A cycle has $\bar{d}=2$, so $M_{2}\geq M$ from Lemma 4.6. Otherwise, we can construct $G$ from a 3-star, a graph with degree sequence $(3,1,1,1)$, by successively adding edges and vertices. For a 3-star, $M_{2}=M=6$. Now, if we add an edge between vertices of degrees $d_{i}$ and $d_{j}$, $M\leftarrow M+2$, and
$M_{2}\leftarrow M_{2}+(d_{i}+1)d_{i}+(d_{j}+1)d_{j}-d_{i}(d_{i}-1)-d_{j}(d_{j}-1)=M_{2}+2(d_{i}+d_{j})\geq M_{2}+2,$
with equality only if we add an edge from a vertex of degree $d_{i}=1$ to a new vertex of degree 0. Otherwise the inequality is strict. The claim now follows by induction. ∎
If $G$ is a path $P_{\ell}$ with $\ell$ vertices, then $M=2(\ell-1)$ and $M_{2}=2(\ell-2)$, so $M-M_{2}=2$. Thus, if $M>M_{2}$, then Lemma 4.7 implies that most components must be paths.
Acknowledgements
The authors thank the anonymous referee for their helpful comments, and thank Brendan McKay for useful discussions. Colin Cooper’s research was supported by DFG Project 491453517. Martin Dyer’s research was supported by an EPSRC grant EP/S016562/1 “Sampling in hereditary classes”. Catherine Greenhill’s research was supported by Australian Research Council Discovery Project DP250101611.